### v1

# LIGHTWEIGHT ERROR MITIGATION STRATEGIES for POST-TRAINING N:M ACTIVATION SPARSITY in LLMS

Abstract
The demand for efficient large language model (LLM) inference has intensified the focus on sparsification techniques. While semi-structured (N:M) pruning is well-established for weights, its application to activation pruning remains underexplored despite its potential for dynamic, input-adaptive compression and reductions in I/O overhead. This work presents a comprehensive analysis of methods for post-training N:M activation pruning in LLMs. Across multiple LLMs, we demonstrate that pruning activations enables superior preservation of generative capabilities compared to weight pruning at equivalent sparsity levels. We evaluate lightweight, plug-and-play error mitigation techniques and pruning criteria, establishing strong hardware-friendly baselines that require minimal calibration. Furthermore, we explore sparsity patterns beyond NVIDIA’s standard 2:4, showing that the 16:32 pattern achieves performance nearly on par with unstructured sparsity. However, considering the trade-off between flexibility and hardware implementation complexity, we focus on the 8:16 pattern as a superior candidate. Our findings provide both effective practical methods for activation pruning and a motivation for future hardware to support more flexible sparsity patterns. Our code is available here.
1 Introduction
The expanding capabilities of Large Language Models (LLMs) have driven enormous demand for efficient AI inference. A common rule of thumb states that inference serving speed for a dense model with $N$ parameters scales as $\propto 1/\sqrt{N}$ (Erdil, 2025). Accelerating inference typically involves either reducing numerical precision via quantization to mitigate memory bottlenecks and enable faster arithmetic (Shvetsov et al., 2024) or sparsification to reduce the number of parameters (Maximov et al., 2025). Sparsity improves LLM efficiency in two key ways: (1) by reducing computation, and (2) by reducing I/O traffic for transferring parameters between memory and compute units – a major bottleneck during inference.
Weights vs. Activations. While sparse weights and activations offer identical theoretical FLOPs (floating point operations count), their practical implications differ. Weight sparsity enables static compression, efficiently reducing model storage and memory bandwidth. However, its static nature often causes irreversible model degradation. In contrast, activation sparsity is dynamic and input-adaptive, preserving model capacity, which we demonstrate in our work.
Sparsity in LLMs. Naturally, pruning is most effective when the values to be pruned already have low magnitudes or are zero, which is often the case for some LLMs’ intermediate representations (Liu et al., 2023), especially due to the ReLU activation function in MLP blocks (Mirzadeh et al., 2023). Even if a model was trained with a non-ReLU activation function, we can still perform relufication Song et al. (2024b) to induce sparsity. This makes activation sparsification a natural choice for enhancing the model’s performance.
Accelerating LLMs with sparse activations. Indeed, recent results demonstrated that activation sparsification supported by specialized software can increase the model’s vector by matrix multiplication during the LLM decoding stage up to $2\timestimes$ (Song et al., 2024b; a; Liu et al., 2024; Lee et al., 2024). The idea behind such approaches is relatively straightforward - if a vector has many zeroes, we can disregard corresponding columns in a weight matrix $W$ on the fly. However, as batch size grows beyond oneone, it becomes challenging to gain benefits (Shrestha et al., 2025). Moreover, most previous methods mainly rely on sparsity induced only by activation functions (Mirzadeh et al., 2023) in MLP blocks.
[…]
Focus and Motivation.
Applying N:M sparsity patterns to activations, while conceptually straightforward, raises several key questions: (Q1) Selection Strategy: As with weight sparsification, the criterion for selecting which activations to prune is crucial for maintaining performance (Zhelnin et al., 2024). (Q2) Error Mitigation: Fine-tuning is a standard method for recovering performance lost to post-training sparsification, but can be prohibitively expensive or risk compromising model safety alignment (Kharinaev et al., 2025). Therefore, we focus on lightweight, plug-and-play and hardware-friendly approaches that require only minimal calibration data (e.g., WikiText) or no data at all to enhance performance. (Q3) Beyond 2:4 Patterns: While 2:4 is the only hardware-supported semi-structured pattern, we explore a broader range {2:4,4:8,8:16,16:32} to motivate future hardware design. While we cannot directly compare acceleration gains between various patterns, our work solely focuses on model performance.
Our Contributions.
This paper makes the following key contributions:
  • Activation vs. Weight Sparsity Superiority: We show that, at similar sparsity levels, activation pruning consistently outperforms weight pruning across four diverse LLMs: Llama2-7B, Llama3-8.1B, Qwen2.5-7B, and Gemma3-4B.
  • Lightweight Error Mitigation and Selection Criteria: We comprehensively evaluate four plug-and-play error mitigation techniques, three evaluated for the first time in the context of activation sparsity. We also evaluate three pruning selection criteria. These methods include statistical criteria such as median shift and variance correction and provide strong, hardware-friendly baselines for activation sparsification.
[…]
2 Hardware and Comparison of 2:4 and 8:16 Sparsity
One of the primary aims of our work is to motivate the design of hardware for more flexible sparsity patterns. Currently, only NVIDIA GPUs support structured sparsity, specifically the 2:4 pattern. This pattern achieves  1.5–1.7x inference acceleration and  1.5x energy reduction for 7B models, primarily from a 2x reduction in memory bandwidth (Fang et al., 2024; Lin et al., 2023). At the same time, any N:M pattern beyond 2:4, including 8:16, can provide the same 2x bandwidth reduction but with greater flexibility. In the general N:MN{:}M format, 2:4 sparsity permits $\binom{4}{2}=6$ non-zero patterns per block, requiring $\log_{2}(6)/4\approx 0.75$ bits per element for metadata. In contrast, the 8:16 pattern allows $\binom{16}{8}=12{,}870$ distinct configurations, demanding $\log_{2}(12{,}870)/16\approx 0.875$ bits per element. Stacking four 2:4 blocks to form a 16-element unit yields only $6^{4}=1{,}296$ configurations—nearly an order of magnitude fewer than the 12,870 offered by a native 8:16 block. Although 8:16 requires slightly more metadata storage, its vastly greater configurational flexibility makes it a compelling candidate. The main trade-off is implementation complexity. Decoding sparsity metadata and gathering non-zero values requires sophisticated circuitry. As the block size $M$ grows, these gather operations could impact performance. However, operating on larger, aligned blocks could also improve cache utilization, potentially amortizing this overhead.
Figure 1: Comparison of unstructured sparsity in activations (ACT) and weights (WT) averaged across four datasets at varying sparsity ratios. Higher is Better. More detailed results are presented in Appendix Table 9.
Figure 2: Comparison of sparsity patterns with unstructured sparsity. 50% and 70% correspond to unstructured sparsity. Lower is Better. More detailed results are presented in Appendix Table 6.
3 Methods
Table 1: Briefly describe evaluated activation pruning metrics (top) and transformations (bottom). Abbreviations in bold with an asterisk (*) denote methods proposed here or first evaluated with sparse activations.
Short Name Method Key Mechanism
Pruning metrics
ACT Magnitude Pruning Selects based on activation magnitude
WT Weight-based Pruning Selects by corresponding weight magnitude
CLACT* Cosine Loss Activation A metric inspired by cosine similarity from Mi et al. (2025)
Amber-Pruner (2025a) Weights Importance A metric which accounts for important weights after outlier removal and normalization
Transformations
D-PTS* Dynamic Per-Token Shift Batch-wise dynamic centering of activations before sparsification
S-PTS (2024) Static Per-Token Shift Fixed centering of activations before sparsification using a per-token bias value pre-collected on WikiText-2
L-PTS* Learnable Per-Token Shift Fixed centering of activations before sparsification using per-token bias value learned on WikiText-2
VAR* Variance Correction Token-wise variance normalization after sparsification
VAR+L-PTS* Scaling + Learnable Shift Apply VAR scaling, then add per-token bias value learned on WikiText-2
R-Sparse (2025) Rank-Aware Sparsity Combines sparse activations with weight low-rank SVD factors learned on WikiText-2
3.1 Preliminaries
Consider a linear layer of a LLM with weight $\mathbf{W}$, input activations $\mathbf{X}$ and output $\mathbf{Y}$ then:
$\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}$ (1)
Our aim is to find a semi-structured pruning mask $\mathbf{M}$ for the activations $\mathbf{X}$ based on some pruning metric $S$, such that:
[…]
where $i,j$ denote the indices of the matrix elements, $\mathrm{t}$ is a threshold value chosen to ensure the desired level of sparsity.
Once the mask $\mathbf{M}$ is computed, the output of the linear layer with pruned activations is given by:
$\mathbf{Y}_{p}=(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}$ (3)
[…]
3.2 Pruning Criterion
In this study, we primarily consider some pruning metrics. You can find brief descriptions of them in 1.
ACT: This is the magnitude activation pruning metric (ACT), defined as the absolute value of the element $\mathbf{X}_{ij}:S_{\textit{ACT}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|$.
WT: This is the weight-based pruning metric (WT), defined as the absolute value of the corresponding weight $\mathbf{W}_{ij}$: $S_{\textit{WT}}(\mathbf{W}_{ij})=|\mathbf{W}_{ij}|$.
CLACT: This metric is inspired by (Mi et al., 2025), which measures the cosine similarity between layer outputs. Building on this concept, we propose a novel metric called Cosine Loss ACTivation (CLACT). This approach approximates the cosine distance between an activation $\mathbf{X}_{ij}$ and its feature vector, quantifying each element’s importance within its activation row. Unlike magnitude-based methods, CLACT identifies activations that are contextually significant rather than merely large in absolute value.
$S_{\textit{CLACT}}(\mathbf{X}_{ij})=\dfrac{|\mathbf{X}_{ij}|}{\sqrt{\sum_{k=1}^{h}\mathbf{X}^{2}_{ik}}}\sqrt{\sum_{p=1}^{l}\mathbf{X}^{2}_{pj}},$ (4)
where $h$ is hidden dimension, $l$ is sequence length, $j$ is a token number and $i$ an element withing this token.
Amber-Pruner: This metric from An et al. (2025b) consists of the following steps: (1) Outlier Removal. To reduce the influence of extreme values, weights $\mathbf{W}$ outside the 0.5th99.5th0.5th-99.5th percentiles are discarded: $\mathbf{W}=\left\{\omega_{k}\mid Q_{0.005}(\mathbf{W})\leq\omega_{k}\leq Q_{0.995}(\mathbf{W})\right\}$. (2) Normalization. The remaining weights are standardized using their mean and variance: $\hat{\mathbf{W}}_{ij}=(\mathbf{W}_{ij}-\mathbb{E}[\mathbf{W}])(\operatorname{Var}[\mathbf{W}])^{-\frac{1}{2}}$. (3) Channel-wise Scoring. For each channel $\hat{\mathbf{W}}_{:,j}$, the $\ell_{2}$ norm is computed, and final scores are assigned to activation elements, $\mathcal{L}(\cdot)$ - denotes channel-wise $\ell_{2}$ norm: $S_{\textit{Amb.-Pr.}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|\cdot\mathcal{L}\left(\hat{\mathbf{W}}_{:,j}\right)$.
Here, we highlight a critical distinction between CLACT and Amber-Pruner. CLACT is a context aware criterion, particularly useful during the prefill stage since with $l=1$, it converges to the $L1$ - norm criterion, moreover, it does not consider weight magnitudes. In contrast, Amber-Pruner explicitly prioritizes weights based on their values and remains effective across both prefill and generation stages but does not consider context.
3.3 Error mitigation and Transformations
Since semi-structured pruning of activations can significantly disrupt the outputs of linear layers, additional transformations can be applied to mitigate these negative effects and restore the pruned activations to their original distribution. You can find brief descriptions of them in 1.
PCS: This approach is inspired by SmoothQuant (Xiao et al., 2023) and has been used for 2:4 weight sparsification (Maximov et al., 2025) as well. It uses a per-channel smoothing (PCS) factor $s$ applied to activations $\mathbf{X}$ before pruning: $\hat{\mathbf{X}}=\mathrm{diag}(s(\mathbf{X},\mathbf{W}))^{-1}\mathbf{X}$. To compensate for this transformation, the weight matrix is correspondingly scaled in the opposite direction. As a result, the output of the linear layer is computed as: $\mathbf{Y}_{p}=(\hat{\mathbf{X}}\odot\mathbf{M})(\mathrm{diag}(s(\mathbf{X},\mathbf{W}))\mathbf{W}^{\top})$. The scale factor $s$ is estimated as Xiao et al. (2023): $s=\sqrt{\max|\mathbf{X}_{:,j}|(\max|\mathbf{W}^{\top}_{:,j}|)^{-1}}$
D-/S-/L-PTS: The next approach applies a dynamic per-token shift (D-PTS) to the input activations $\mathbf{X}$ using a bias term $\mathbf{\eta}$, such that $\hat{\mathbf{X}}=\mathbf{X}-\mathbf{\eta}$. The purpose of this shift is to center most elements of the activations near zero. As a compensatory transformation, the bias $\eta$ is added to the weights, so the output of the linear layer is given by Chua et al. (2024): $\mathbf{Y}_{p}=((\hat{\mathbf{X}}\odot\mathbf{M})+\mathbf{\eta})\mathbf{W}^{\top}$, where $\eta=\overline{\mathbf{X}}$. A related variant is the static per-token shift (S-PTS), where the bias $\eta$ is collected during a short warm-up phase and then fixed for all subsequent batches. This reduces computational overhead while providing similar centering effects. Finally, we introduce a learnable per-token shift (L-PTS).
VAR: In contrast to the previous approach, this method applies a per-token scaling factor $\mathbf{\nu}$ to pruned activations such that the output matrix is given as: $\mathbf{Y}_{p}=\mathbf{\nu}(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}$, The scale factor $\mathbf{\nu}$ is determined as
$\nu=\sqrt{\dfrac{\operatorname{Var}[\mathbf{X}]}{\operatorname{Var}[\mathbf{X}\odot\mathbf{M}]}}$ (5)
We refer to this method as variance correction VAR.
VAR+L-PTS: This method is a combination of the per-token scaling with factor VAR $\mathbf{\nu}$ is given by Eq. 5 and the shift L-PTS.
R-Sparse: This method from  (Kamirul et al., 2025) combines activation sparsity with a low-rank approximation of the weight matrix. We provide more details in Appendix A.
Throughout our experiments we use the WikiText-2 dataset for calibration and learning some parameters. For S-PTS we calibrate a single per-channel bias vector $\eta$ and then keep it fixed for all subsequent batches. For the learnable variant L-PTS (and the combined VAR+L-PTS) we train $\eta$ and fine-tune other weights. For R-Sparse we train in the same way the low-rank factors of the weight adaptation. In all other methods no parameters are learned or calibrated.
3.4 Evaluation & Models
We begin with a broad funnel of experiments to evaluate all methods listed in Table 1. At this initial stage, we use the Core Datasets, Single-Choice and Multi-Choice benchmarks including BoolQ, WinoGrande, PIQA, and ARC-Easy. Once this exploratory phase is complete, we narrow our focus to the most promising approaches, evaluating them primarily on two models: Qwen2.5-7B-Instruct and Llama3-8B-Instruct. This more targeted evaluation phase not only refines the set of methods under consideration but also expands the evaluation to the Extended Datasets: HellaSwag, OpenBookQA, RTE, MMLU, Lambada_standard, Lambada_openai, and IFEval. For calibration tasks, when required, we use WikiText-2. This dataset selection follows established model compression practices Egiazarian et al. (2024); Frantar et al. (2022); van Baalen et al. (2024). All evaluations are performed using the LM Eval Harness Gao et al. (2023), and full dataset details are provided in Table 8. For the models we focus on: Qwen2.5-7B-Instruct, Llama3-8B-Instruct, Llama2-7B, Gemma3-4B-Instruct. Importantly, for Qwen2.5-7B-Instruct, we did not sparsify the key, query, or value activations, as preliminary experiments showed severe model degradation when these layers were pruned.
4 Results
4.1 Sparse Weights vs. Sparse Activations
In Figure 2 and Table 9, we demonstrate that unstructured weight sparsification causes greater model degradation than unstructured activation sparsification at the same sparsity levels: {20%, 50%, 70%, 90%}. For this evaluation, we specifically use unstructured magnitude-based sparsification, as it is less damaging than semi-structured sparsification and thus provides a lower bound on performance degradation.
4.2 Optimal semi-structured sparsity patterns
Our preliminary investigation demonstrates that while the 16:32 pattern achieves performance close to unstructured sparsity (a 5.4% drop versus 4.5% for 50% unstructured), it requires more metadata and greater resources for gather operations, as discussed in Section 2. Therefore, we focus on the 8:16 pattern, despite its higher performance drop of 7.38%. For comparison, the 2:4 pattern results in a 14.35% drop. These results are shown in Figure 2 and Table 6 in Appendix, we used only magnitude pruning to obtain these results. By demonstrating the superior model quality of 8:16 sparsity, our work incentivizes hardware designers to invest in—or at least consider—supporting the 8:16 pattern.
Table 2: Average relative performance (%) across four datasets. Values indicate performance drops (lower is better), negative values signify performance improvement. Methods marked with an asterisk (*) are proposed in this paper. Full, non-aggregated results are available in Appendix 10.
Models
Method Llama2-7B Qwen2.5-7B Gemma3-4B LLama3-8B Average Drop ($\downarrow)$
Activations Unstructured 50%
ACT 2.312.31% 3.873.87% 4.804.80% 4.304.30% 3.823.82%
2:4 Weight Sparsity
WT 16.5216.52% 12.9612.96% 34.8634.86% 33.6333.63% 24.4924.49%
2:4 Activations Selection Criteria
ACT 9.439.43% 4.954.95% 9.949.94% 14.3514.35% 9.679.67%
CLACT* 8.32% 2.45-2.45% 8.018.01% 17.2717.27% 7.79%
Amber-Pruner 11.7011.70% 1.23-1.23% 5.915.91% 15.0115.01% 7.857.85%
2:4 Activations Transformations
VAR* 9.769.76% 1.48-1.48% 2.96% 13.1113.11% 6.096.09%
D-PTS* 10.6710.67% -6.46% 4.584.58% 14.5914.59% 5.845.84%
S-PTS 10.3710.37% 4.43-4.43% 3.933.93% 7.31% 4.29%
L-PTS* 13.1313.13% 3.663.66% 4.214.21% 14.1314.13% 8.798.79%
R-SPARSE (64) 12.9012.90% 2.55-2.55% 5.175.17% 15.2815.28% 7.707.70%
R-SPARSE (128) 12.2312.23% 1.51-1.51% 5.165.16% 16.3416.34% 8.058.05%
8:16 Weight Sparsity
WT 7.847.84% 9.549.54% 26.1126.11% 27.2627.26% 17.6817.68%
8:16 Activations Selection Criteria
ACT 5.375.37% 4.384.38% 4.764.76% 7.387.38% 5.475.47%
CLACT* 3.983.98% 4.02-4.02% 0.600.60% 8.608.60% 2.292.29%
Amber-Pruner 5.325.32% 6.28-6.28% 0.080.08% 7.137.13% 1.56%
8:16 Activations Transformations
VAR* 4.854.85% 1.931.93% -1.87% 8.308.30% 3.303.30%
D-PTS* 4.634.63% -8.28% 5.165.16% 6.79% 2.072.07%
S-PTS 3.87% 7.24-7.24% 1.54-1.54% 7.37.3% 0.61%
L-PTS* 8.158.15% 1.711.71% 4.214.21% 7.197.19% 5.325.32%
R-SPARSE (64) 5.915.91% 6.91-6.91% 1.36-1.36% 8.448.44% 1.521.52%
R-SPARSE (128) 7.937.93% 5.44-5.44% 0.44-0.44% 8.448.44% 2.632.63%
4.3 Evaluation of Pruning Selection Criteria on Single/Multi-choice Datasets
We evaluate CLACT, Amber-Pruner, and magnitude pruning as a baseline. The main results for the 2:4 and 8:16 sparsity patterns are presented in Table 2. On average, both CLACT and Amber-Pruner outperform magnitude pruning by at least 2%, however, we observe no clear winner between them. As noted in Section 3.2, these criteria are designed for different purposes: CLACT adjusts based on context, while Amber-Pruner adjusts based on weight magnitudes. Notably, for Llama3-8B under the 2:4 sparsity pattern, simple magnitude pruning outperforms both advanced criteria, underscoring model and architecture-specific sensitivities to pruning strategies.
4.4 Evaluation of Transformations on Single/Multi-choice Datasets
Our main results are presented in Table 2. Surprisingly, we find that simple methods such as dynamic and static per-token shifts (D-PTS, S-PTS) outperform most other approaches. The second most effective methods are VAR and R-SPARSE. We also observe that increasing the number of dimensions in R-SPARSE (from 64 to 128) leads to worse performance, which may indicate overfitting on the calibration data. Finally, we note that L-PTS, the approach with learnable per-token shifts, significantly underperforms compared to its static counterpart, S-PTS.
4.5 Evaluation transformations with Instruction-Following Tasks
Table 3 presents instruction-following performance on the IFEval benchmark for Llama3-8B and Qwen2.5-7B, evaluated under two semi-structured sparsity patterns (2:4 and 8:16) and four activation transformation methods: S-PTS, D-PTS, R-Sparse, and VAR. First of all, we observe a strong model degradation on generative tasks. Second of all, we see that VAR is the strongest performer overall, especially for Llama3-8B. S-PTS/D-PTS are competitive and lightweight, and R-Sparse lags significantly, particularly at 2:4. We speculate that while semi-structured patterns are good for prefill stage in LLMs they significantly degrade performance during decode stage. However, as discussed in Section 1 decode stage for single vector can be accelerated with more flexible approaches.
Table 3: Evaluation with Instruction-Following (IFEval) for Llama3-8B and Qwen2.5-7B. PS denotes prompt-level strict acc, PL denotes prompt-level loose acc.
Model Original Pattern S-PTS D-PTS R-Sparse VAR
PS PL PS PL PS PL PS PL PS PL
Llama3-8B 0.4455 0.4861 2:4 0.1682 0.1904 0.1941 0.2015 0.0869 0.0979 0.2237 0.2458
8:16 0.2995 0.3327 0.2828 0.3198 0.2089 0.2311 0.3161 0.3586
Qwen2.5-7B 0.7135 0.7394 2:4 0.4325 0.5176 0.4399 0.5120 0.2736 0.3457 0.4565 0.5342
8:16 0.5194 0.5804 0.5434 0.5989 0.3697 0.4196 0.5249 0.5896
4.6 Selection Criteria and Transformations for Unstructured Pruning
Although unstructured pruning offers limited efficiency gains compared to structured variants, it serves as a valuable proxy task for evaluating the robustness and generalization of proposed methods. We evaluate D-PTS, VAR, and two selection criteria in this experiment: CLACT and Amber-Pruner using the Llama3-8B model. Results in Table 4 indicate that VAR is the most effective transformation under unstructured sparsity. Moreover, CLACT outperforms Amber-Pruner by a wider margin here than in our semi-structured pruning experiments. These findings suggest two key insights: (1) No single method emerges as optimal for both unstructured and semi-structured sparsity. (2) The methods proposed in this work, VAR and CLACT, demonstrate strong generalization and are well-suited for unstructured activation pruning.
Table 4: Performance comparison of pruning methods under 50% and 70% unstructured sparsity for Llama3-8B model.
Model ARC Easy BoolQ PIQA WinoGrande Avg. Drop (%)
Original 0.8207 0.8391 0.8003 0.7340
Unstructured 50%
ACT 0.7770 0.8198 0.7714 0.6858 4.45
D-PTS 0.78577 0.8253 0.7807 0.68981 3.60
VAR 0.78409 0.8192 0.77584 0.7048 3.47
CLACT 0.7803 0.8253 0.7655 0.70007 3.89
Amber-Pruner 0.768 0.82018 0.7627 0.7016 4.45
Unstructured 70%
ACT 0.55808 0.63119 0.6474 0.5477 25.32
D-PTS 0.56481 0.62415 0.6507 0.53433 25.68
VAR 0.61447 0.65107 0.67573 0.5319 22.66
CLACT 0.5551 0.6039 0.62676 0.52407 27.67
Amber-Pruner 0.48737 0.5938 0.5897 0.539 30.68
4.7 Combination of Transformations and Pruning criteria
Next, we evaluated combinations of multiple approaches to explore potential performance gains. These combinations and their results are presented in Table 7. As shown, none of the evaluated combinations outperforms any single method, highlighting the challenges of naively combining them.
Table 5: Aggregated results on Llama3-8B with 8:16 activation sparsity with different layers pruned. ORIG. AVG. denotes original model performance. Layers indicates the subset of linear layers where the method was applied. Non aggregated results are presented in Appendix 12. Drop is computed without accounting for perplexity.
ORIG. AVG. Method Layers PPL AVG. Drop $\downarrow$
LS+L-PTS all 9.6036 0.6047 10.90%
LS+L-PTS key,out,gate,down 8.3483 0.6385 5.43%
0.6726 LS+L-PTS key,value,gate,down 8.0821 0.6503 3.56%
LS+L-PTS + VAR all 9.4983 0.6056 10.60%
LS+L-PTS + VAR key,out,gate,down 8.2930 0.6422 4.64%
LS+L-PTS + VAR key,value,gate,down 8.0259 0.6516 3.36%
4.8 Layer Sensitivity
The next question we investigate in this work is layer sensitivity to activation sparsity. For this experiment, we use the Extended Datasets, the Llama3-8B model, and the 8:16 sparsity pattern. The main results are presented in Table 5. Here, we focus on methods with learnable parameters, primarily L-PTS and LS (a learnable, diagonal scaling projection). While in most of our experiments, learnable methods underperformed on average compared to static or magnitude-based approaches, the best-performing configurations for Llama3-8B under 8:16 sparsity were, in fact, those employing learnable parameters — as shown in Table 5. We find that activations from the up projection (in the FFN) and the out projection [^1] The output projection of the attention mechanism. It combines outputs from all attention heads and projects them back to the model’s hidden dimension. are the most sensitive to sparsification, pruning these layers leads to the most significant performance drops, suggesting they should be preserved or treated with greater care in compression strategies. While we can not say this generalizes across all layers, we empirically demonstrate that some layers are more important than others.
5 Discussion
The performance gap between multiple/single-choice benchmarks (e.g., BoolQ, PIQA) and IFEval likely stems from differences in the inference stages they emphasize. Core QA benchmarks primarily stress the prefill phase, whereas IFEval evaluates both prefill and autoregressive generation. Our evaluation remains valid: semi-structured patterns like 2:4 and 8:16 are especially effective at accelerating the prefill stage, which often dominates inference latency.
6 Limitations
Our work has three main limitations. First, evaluations rely on software emulation, precluding real-world measurements of speedup or energy savings on hardware supporting sparsity beyond 2:4. Second, our layer sensitivity analysis is preliminary, while up-projection and out-projection layers appear most vulnerable, broader architectural studies are needed. Third, Qwen2.5-7B shows anomalous gains on multiple-choice benchmarks under certain sparsification methods, an artifact we attribute to benchmark limitations, as no such gains appear on IFEval. Critically, generative performance degrades significantly, underscoring the need for evaluation beyond multiple-choice QA.
7 Conclusion
This work establishes that post-training activation pruning is more accuracy preserving than weight pruning for Large Language Models. Across four diverse LLMs, we demonstrate that activation pruning consistently preserves a model’s capabilities better than weight pruning at equivalent sparsity levels, highlighting its potential for dynamic, input-adaptive efficiency gains.
Our key novel contributions are twofold. First, we conduct the first comprehensive evaluation of lightweight, plug-and-play error mitigation techniques for activation sparsity, including the introduction and evaluation of three new methods: Cosine Loss Activation (CLACT) as a context-aware pruning criterion, Dynamic/Learnable Per-Token Shift (D-PTS/L-PTS) and Variance Correction (VAR). We show that these simple methods, often outperforming more complex approaches, providing strong, hardware-friendly baselines for the community. Second, we systematically explore semi-structured sparsity patterns beyond the hardware standard 2:4. We find that the 8:16 pattern offers more flexibility and, ultimately, lower model degradation. Although the 16:32 pattern performs even closer to unstructured sparsity, we advocate for 8:16 as the optimal target for future hardware design due to its balance of performance gain and implementation feasibility. In summary, this work not only provides practical, high-performing methods for post-training activation pruning but also creates grounded motivation for 8:16 structured sparsity support, which would unlock significant efficiency gains for LLM inference with minimal added complexity.
Acknowledgment on LLM assisted writing:
This paper used open access Qwen3-Max, in some parts of the paper, for proofreading and text rephrasing in accordance with formal style.
Appendix
Appendix A R-Sparse Details
Finally, we include R-Sparse (Kamirul et al., 2025), which combines activation sparsity with a low-rank approximation of the weight matrix. Instead of pruning solely by magnitude, R-Sparse decomposes the computation into two parts: (i) sparse channels with high-magnitude activations, and (ii) a low-rank component obtained via SVD of $\mathbf{W}$ that approximates the contribution of pruned activations.
Formally, the linear layer
$\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}$ (6)
is approximated as
$\mathbf{Y}\approx\mathbf{Y}_{s}+\mathbf{Y}_{r},$ (7)
where
$\mathbf{Y}_{s}=\sigma_{t(s)}(\mathbf{X})\mathbf{W}^{\top},\quad\mathbf{Y}_{r}=(\mathbf{X}-\sigma_{t(s)}(\mathbf{X}))(A_{r}B_{r})^{\top}.$ (8)
Here $\sigma_{t(s)}(\cdot)$ denotes sparsification of activations with threshold $t(s)$, and $A_{r}B_{r}^{\top}$ is the rank-$r$ approximation of $\mathbf{W}$ obtained from its truncated SVD. The trade-off between $\mathbf{Y}_{s}$ and $\mathbf{Y}_{r}$ is determined by a sparsity budget $s$ and rank $r$, which can be optimized via evolutionary search.
Appendix B Different patterns for semi-structure sparsification
Table 6: Performance comparison of different sparsity patterns on Llama3-8B across various benchmarks. Values represent accuracy scores, with the last column showing the average performance drop relative to the original model.
ARC Easy BoolQ PIQA WinoGrande Avg Drop ($\downarrow$)
Original $0.8207$ $0.8391$ $0.8003$ $0.7340$
2:4 $0.6837$ $0.7261$ $0.7163$ $0.6110$ 14.3514.35%
4:8 $0.7272$ $0.7810$ $0.7529$ $0.6393$ 9.299.29%
8:16 $0.7525$ $0.7969$ $0.7568$ $0.6551$ 7.387.38%
16:32 $0.7698$ $0.8082$ $0.7688$ $0.6771$ 5.405.40%
50% unstructured $0.7820$ $0.8198$ $0.7714$ $0.6858$ 4.304.30%
70% unstructured $0.5580$ $0.6311$ $0.6474$ $0.5477$ 25.3225.32%
B.1 Comparison of different patterns
Table 7: A comparison of combined approaches with 8:16 semi-structured sparsity. Average relative performance (%) across four datasets. Values indicate performance drops (lower is better), negative values signify performance improvement. Full, non-aggregated results are available in Appendix 10.
Models
Method Llama2-7B Qwen2.5-7B Gemma3-4B LLama3-8B Average Drop ($\downarrow)$
CLACT + PTS 5.635.63% 5.06-5.06% 0.500.50% 8.558.55% 2.40%
CLACT + VAR 5.075.07% 2.90-2.90% 0.540.54% 8.598.59% 2.822.82%
Amber-Pruner + PTS 6.166.16% 3.47-3.47% 0.170.17% 7.427.42% 2.572.57%
Amber-Pruner + VAR 4.744.74% 3.63-3.63% 0.16-0.16% 8.398.39% 2.34%
L-PTS + VAR 6.876.87% 2.862.86% 3.413.41% 7.157.15% 5.075.07%
Appendix C Datasets
Table 8: Datasets used to evaluate hypotheses. Prompt-level strict accuracy is the fraction of prompts for which all verifiable instructions in the prompt are followed exactly as stated. Instruction-level strict accuracy is the fraction of individual instructions that are followed exactly as stated, averaged across all instructions.
Dataset Description Metric
WikiText-2 (Merity et al., 2016) A collection of over 100 million tokens extracted from the set of verified Good and Featured articles on Wikipedia. Perplexity
ARC-Easy (Clark et al., 2018) QA benchmark for genuine grade-school level, multiple-choice science questions. The dataset contains 2251 examples for training, 570 for development and 2376 for testing. Accuracy
ARC_Challenge (Clark et al., 2018) QA benchmark for more difficult grade-school level science questions, part of the AI2 Reasoning Challenge. Designed to require deeper reasoning than ARC-Easy. Accuracy
BoolQ (Clark et al., 2019) QA benchmark for yes/no questions. The dataset contains 9427 examples for training and 3270 for testing. Accuracy
PIQA (Bisk et al., 2020) Physical commonsense QA benchmark for choosing the right answer between two options. Contains 16K train, 2K dev, and 3K test examples. Accuracy
WinoGrande (Sakaguchi et al., 2020) QA benchmark for pronoun resolution with adversarial filtering. Contains 40K train, 1267 dev, and 1767 test examples. Accuracy
HellaSwag (Zellers et al., 2019) Commonsense reasoning benchmark for sentence completion, designed to be easy for humans but hard for models. Contains 70K train and 10K validation examples. Accuracy
OpenBookQA (Mihaylov et al., 2018) Open-book question answering dataset requiring retrieval of elementary science facts. Contains 5957 4-way multiple-choice questions. Accuracy
RTE (Dagan et al., 2005; Bar-Haim et al., 2006) Recognizing Textual Entailment datasets from PASCAL challenges. Task is to classify if a hypothesis is entailed by a premise. Accuracy
MMLU (Hendrycks et al., 2020) Massive Multitask Language Understanding benchmark covering 57 subjects across STEM, humanities, and social sciences. Measures multitask accuracy. Accuracy
Lambada_Standard (Paperno et al., 2016) Word prediction task requiring broad discourse context. Target word is unpredictable from local context alone. Accuracy
Lambada_OpenAI (Paperno et al., 2016) LAMBADA test set preprocessed by OpenAI for standardized evaluation. Task remains final word prediction with long-range dependencies. Accuracy
GSM8K (Cobbe et al., 2021) Grade school math word problems requiring multi-step reasoning. Contains 7.5K train and 1.3K test examples. Accuracy (Strict) Accuracy (Flexible)
IFEval (Zhou et al., 2023) Benchmark with 541 prompts containing verifiable instructions to measure instruction-following fidelity. Accuracy (Prompt-level) Accuracy (Instruct-level)
Appendix D Weights versus Activations
Table 9: The performance of models with applied unstructured activation pruning. We show that even with severe sparsity (70-90%) models were able to perform decently on our benchmarks. ACT stands for activations pruning, WT — for weight pruning. OUT denotes values more than $2^{3}$, according accuracy scores most likely correspond to random.
Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Drop $(\downarrow)\%$
Llama2-7B
   Base 6.94 0.74 0.80 0.76 0.66 -
   0.2 ACT 6.96 0.74 0.80 0.77 0.66 -$\textbf{-0.33}\%$
   0.2 WT 7.49 0.72 0.80 0.76 0.66 $0.68\%$
   0.5 ACT 7.53 0.70 0.78 0.75 0.66 $\textbf{2.32}\%$
   0.5 WT 18.72 0.60 0.72 0.70 0.61 $11.10\%$
   0.7 ACT 20.11 0.56 0.64 0.65 0.53 $\textbf{19.62}\%$
   0.7 WT OUT 0.27 0.38 0.54 0.47 $43.44\%$
   0.9 ACT OUT 0.26 0.38 0.52 0.49 $43.39\%$
   0.9 WT OUT 0.27 0.38 0.53 0.48 $43.39\%$
Qwen2.5-7B
   Base 7.46 0.69 0.86 0.75 0.60 -
   0.2 ACT 7.48 0.69 0.86 0.74 0.61 $\textbf{2.37}\%$
   0.2 WT 8.03 0.67 0.86 0.74 0.60 $3.42\%$
   0.5 ACT 8.3 0.67 0.87 0.74 0.58 $3.87\%$
   0.5 WT 43.6 0.56 0.80 0.68 0.57 $\textbf{3.42}\%$
   0.7 ACT 18.7 0.6 0.81 0.70 0.58 $\textbf{3.87}\%$
   0.7 WT OUT 0.28 0.38 0.54 0.48 $12.12\%$
   0.9 ACT OUT 0.25 0.38 0.54 0.52 $44.22\%$
   0.9 WT OUT 0.25 0.58 0.54 0.51 $\textbf{36.35}\%$
Gemma3-4B
   Base 17.29 0.72 0.84 0.72 0.62 -
   0.2 ACT 17.60 0.71 0.84 0.72 0.60 $\textbf{3.35}\%$
   0.2 WT 18.93 0.68 0.84 0.72 0.59 $4.74\%$
   0.5 ACT 22.39 0.71 0.83 0.72 0.57 $\textbf{4.80}\%$
   0.5 WT 273 0.36 0.55 0.61 0.52 $30.89\%$
   0.7 ACT 88 0.55 0.63 0.66 0.54 $\textbf{19.57}\%$
   0.7 WT OUT 0.27 0.49 0.53 0.51 $38.81\%$
   0.9 ACT OUT 0.26 0.38 0.54 0.50 $42.64\%$
   0.9 WT OUT 0.25 0.45 0.52 0.52 $\textbf{40.60}\%$
Appendix E Main extended results
In this section we present non - aggregated results for all of the methods and datasets., results are presented in Table 11 and Table 10.
Table 10: Semi-Structured 2:4 Sparsification - performance Metrics, for calibration, when it is required, and perplexity we use WikiText2. Average Drop is computed without accounting for perplexity
Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Average Drop %
   Llama2-7B 6.94 0.74 0.80 0.76 0.66 -
ACT 10.23 0.66 0.71 0.71 0.60 9.43%
WT 42.40 0.57 0.65 0.69 0.56 16.52%
D-PTS 9.38 0.64 0.68 0.71 0.61 10.67%
S-PTS 9.36 0.66 0.68 0.71 0.60 10.37%
VAR 8.31 0.67 0.69 0.72 0.59 9.76%
CLACT 8.23 0.65 0.72 0.71 0.63 8.32%
Amber-Pruner 9.24 0.64 0.68 0.69 0.60 11.70%
LPTS 8.89 0.65 0.60 0.72 0.59 13.13%
LPTS + VAR 8.39 0.67 0.63 0.72 0.60 11.47%
R-SPARSE (64) 9.19 0.66 0.63 0.69 0.59 12.90%
R-SPARSE (128) 9.29 0.65 0.65 0.70 0.59 12.23%
   Llama3.1-8B 7.21 0.82 0.84 0.80 0.73 -
ACT 16.61 0.68 0.73 0.72 0.61 14.35%
WT 20.14 0.41 0.57 0.60 0.54 33.63%
PTS 16.4 0.69 0.73 0.72 0.60 14.59%
S-PTS (N-100) 16.5 0.67 0.74 0.72 0.60 14.61%
S-PTS (N-200) 16.5 0.68 0.73 0.72 0.61 14.31%
VAR 14.17 0.70 0.73 0.73 0.62 13.11%
CLACT 19.49 0.65 0.71 0.69 0.59 17.27%
WANDA 15.86 0.66 0.74 0.69 0.61 15.01%
L-PTS 12.77 0.71 0.71 0.73 0.59 14.13%
L-PTS + VAR 12.40 0.73 0.71 0.73 0.60 13.49%
R-SPARSE (64) 15.07 0.69 0.72 0.71 0.61 15.28%
R-SPARSE (128) 16.09 0.67 0.71 0.70 0.61 16.34%
   Qwen2.5-7B 7.46 0.69 0.86 0.75 0.60 -
ACT 10.06 0.65 0.86 0.72 0.54 4.95%
WT 35.37 0.53 0.78 0.68 0.54 12.96%
D-PTS 10.07 0.79 0.86 0.76 0.66 -6.46%
S-PTS 10.74 0.78 0.84 0.74 0.65 -4.43%
VAR 13.95 0.74 0.83 0.74 0.61 -1.48%
CLACT 11.16 0.73 0.84 0.71 0.67 -2.45%
Amber-Pruner 10.64 0.74 0.84 0.70 0.64 -1.23%
LPTS 9.13 0.67 0.81 0.72 0.58 3.66%
LPTS + VAR 9.10 0.68 0.81 0.73 0.56 3.97%
R-SPARSE (64) 9.03 0.79 0.76 0.75 0.64 -2.55%
R-SPARSE (128) 9.12 0.77 0.77 0.75 0.63 -1.51%
   Gemma3-4B 17.29 0.72 0.84 0.72 0.62 -
ACT 35.62 0.65 0.76 0.70 0.51 9.94%
WT 421.95 0.35 0.44 0.58 0.49 34.86%
D-PTS 35.94 0.70 0.76 0.70 0.60 4.58%
S-PTS 35.84 0.71 0.77 0.70 0.60 3.93%
VAR 33.25 0.60 0.76 0.63 0.54 5.04%
CLACT 39.22 0.66 0.74 0.67 0.59 8.01%
Amber-Pruner 35.56 0.67 0.76 0.68 0.61 5.91%
LPTS 19.55 0.65 0.73 0.70 0.55 9.19%
LPTS + VAR 19.13 0.65 0.74 0.71 0.53 9.82%
R-SPARSE (64) 17.04 0.69 0.76 0.69 0.60 5.17%
R-SPARSE (128) 16.17 0.68 0.75 0.70 0.61 5.16%
Table 11: Semi-Structured 8:16 Sparsification - performance Metrics, for calibration, when it is required, and perplexity we use WikiText2. Average Drop is computed without accounting for perplexity.
Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Average Drop %
   Llama2-7B 6.94 0.74 0.80 0.76 0.66 -
ACT 8.12 0.69 0.75 0.73 0.63 5.37%
WT 20.47 0.64 0.76 0.72 0.61 7.84%
D-PTS 6.92 0.70 0.73 0.75 0.64 4.63%
S-PTS 6.93 0.70 0.73 0.75 0.66 3.87%
VAR 6.67 0.69 0.72 0.75 0.65 4.85%
CLACT 6.54 0.71 0.74 0.75 0.64 3.98%
CLACT + PTS 7.00 0.69 0.72 0.73 0.64 5.63%
CLACT + VAR 6.72 0.69 0.73 0.75 0.64 5.07%
R-SPARSE (64) 7.75 0.69 0.71 0.73 0.64 5.91%
R-SPARSE (128) 7.82 0.68 0.69 0.74 0.61 7.93%
Amber-Pruner 8.10 0.66 0.75 0.73 0.66 5.32%
Amber-Pruner + PTS 6.90 0.68 0.72 0.72 0.65 6.16%
Amber-Pruner + VAR 6.66 0.70 0.72 0.74 0.65 4.74%
LPTS 7.50 0.69 0.66 0.74 0.63 8.15%
LPTS + VAR 7.52 0.69 0.67 0.74 0.64 6.87%
   Llama3.1-8B 7.21 0.82 0.84 0.80 0.73 -
ACT 10.32 0.75 0.80 0.76 0.66 7.38%
WT 22.56 0.51 0.64 0.63 0.54 27.26%
D-PTS 10.34 0.76 0.80 0.76 0.66 6.79%
S-PTS 10.31 0.76 0.80 0.75 0.66 7.30%
VAR 10.67 0.74 0.79 0.75 0.66 8.30%
CLACT 10.67 0.73 0.79 0.74 0.66 8.60%
CLACT + PTS 10.68 0.74 0.79 0.74 0.65 8.55%
CLACT + VAR 10.15 0.74 0.79 0.75 0.64 8.59%
R-SPARSE (64) 11.42 0.75 0.77 0.75 0.66 8.44%
R-SPARSE (128) 10.43 0.75 0.78 0.74 0.66 8.49%
Amber-Pruner 10.16 0.73 0.80 0.75 0.68 7.13%
Amber-Pruner + PTS 10.17 0.75 0.80 0.75 0.66 7.42%
Amber-Pruner + VAR 9.94 0.74 0.80 0.75 0.64 8.39%
LPTS 10.04 0.76 0.79 0.77 0.65 7.19%
LPTS + VAR 10.26 0.77 0.78 0.76 0.66 7.15%
   Qwen2.5-7B 7.46 0.69 0.86 0.75 0.60 -
ACT 8.61 0.66 0.87 0.73 0.53 4.38%
WT 40.79 0.59 0.82 0.67 0.52 9.54%
D-PTS 8.61 0.80 0.87 0.77 0.68 -8.28%
S-PTS 8.84 0.80 0.86 0.76 0.67 -7.24%
VAR 11.91 0.69 0.72 0.75 0.65 1.93%
CLACT 8.94 0.77 0.85 0.73 0.65 -4.02%
CLACT + PTS 8.94 0.77 0.86 0.83 0.67 -5.06%
CLACT + VAR 8.87 0.76 0.84 0.71 0.65 -2.90%
R-SPARSE (64) 8.12 0.82 0.79 0.77 0.69 -6.90%
R-SPARSE (128) 8.24 0.80 0.79 0.77 0.67 -5.40%
Amber-Pruner 8.80 0.77 0.86 0.74 0.69 -6.20%
Amber-Pruner + PTS 8.79 0.77 0.85 0.73 0.64 -3.40%
Amber-Pruner + VAR 8.73 0.75 0.85 0.74 0.65 -3.60%
LPTS 8.23 0.69 0.83 0.75 0.57 1.70%
LPTS + VAR 8.21 0.70 0.83 0.73 0.56 2.80%
   Gemma3-4B 17.29 0.72 0.84 0.72 0.62 -
ACT 25.31 0.70 0.81 0.71 0.55 4.76%
WT 198.53 0.39 0.60 0.62 0.52 26.11%
D-PTS 25.17 0.70 0.81 0.71 0.54 5.16%
S-PTS 25.40 0.75 0.82 0.74 0.63 -1.54%
VAR 23.93 0.75 0.81 0.73 0.65 -1.87%
CLACT 25.85 0.75 0.81 0.71 0.61 0.60%
CLACT + PTS 26.03 0.73 0.81 0.70 0.63 0.50%
CLACT + VAR 24.78 0.74 0.81 0.70 0.63 0.54%
R-SPARSE (64) 15.39 0.76 0.80 0.74 0.64 -1.36%
R-SPARSE (128) 14.55 0.74 0.80 0.73 0.63 -0.44%
Amber-Pruner 25.11 0.74 0.82 0.70 0.63 0.08%
Amber-Pruner + PTS 25.28 0.75 0.81 0.70 0.63 0.17%
Amber-Pruner + VAR 23.97 0.74 0.81 0.71 0.64 -0.16%
LPTS 15.73 0.70 0.79 0.73 0.56 4.21%
LPTS + VAR 15.68 0.71 0.79 0.72 0.57 3.41%
Table 12: Llama3-8B with 8:16 activation sparsity. LS+L-PTS indicates Learnable Diagonal Scale + Learnable Shift, “Layers” indicates the subset of linear layers where the method was applied. Drop is computed without accounting for perplexity.
Method Layers PPL BoolQ WinoGrande PIQA ARC Easy ARC Chal. HellaSwag OpenBookQA RTE MMLU Lambada stand. Lambada (OpenAI) Drop % $\downarrow$
ORIGINAL 0.8391 0.7340 0.8003 0.8207 0.5196 0.5905 0.3420 0.6859 0.6790 0.6569 0.7308
LS+L-PTS all 9.6036 0.7841 0.6638 0.7715 0.7647 0.4514 0.5294 0.2720 0.6318 0.5521 0.5686 0.6625 10.90%
LS+L-PTS k,o,gate,down 8.3483 0.8205 0.7111 0.7889 0.7887 0.4659 0.5591 0.3200 0.6462 0.6060 0.6123 0.7046 5.43%
LS+L-PTS k,v,gate,down 8.0821 0.8352 0.7174 0.7867 0.7992 0.4898 0.5651 0.3260 0.6643 0.6322 0.6262 0.7108 3.56%
LS+L-PTS + VAR all 9.4983 0.7872 0.6606 0.7606 0.7601 0.4334 0.5372 0.2880 0.6390 0.5532 0.5729 0.6689 10.60%
LS+L-PTS + VAR k,o,gate,down 8.2930 0.8116 0.7135 0.7851 0.7908 0.4838 0.5634 0.3300 0.6498 0.6095 0.6189 0.7079 4.64%
LS+L-PTS + VAR k,v,gate,down 8.0259 0.8306 0.7269 0.7851 0.7955 0.4863 0.5673 0.3260 0.6715 0.6327 0.6317 0.7143 3.36%

### v2

@@
-LIGHTWEIGHT ERROR MITIGATION STRATEGIES for POST-TRAINING N:M ACTIVATION SPARSITY in LLMS
+Motivating Next-Gen Accelerators with Flexible $N:M$ Activation Sparsity via Benchmarking Lightweight Post-Training Sparsification Approaches
 Abstract
-The demand for efficient large language model (LLM) inference has intensified the focus on sparsification techniques.
-While semi-structured (N:M) pruning is well-established for weights, its application to activation pruning remains underexplored despite its potential for dynamic, input-adaptive compression and reductions in I/O overhead.
-This work presents a comprehensive analysis of methods for post-training N:M activation pruning in LLMs.
-Across multiple LLMs, we demonstrate that pruning activations enables superior preservation of generative capabilities compared to weight pruning at equivalent sparsity levels.
-We evaluate lightweight, plug-and-play error mitigation techniques and pruning criteria, establishing strong hardware-friendly baselines that require minimal calibration.
-Furthermore, we explore sparsity patterns beyond NVIDIA’s standard 2:4, showing that the 16:32 pattern achieves performance nearly on par with unstructured sparsity.
-However, considering the trade-off between flexibility and hardware implementation complexity, we focus on the 8:16 pattern as a superior candidate.
-Our findings provide both effective practical methods for activation pruning and a motivation for future hardware to support more flexible sparsity patterns.
-Our code is available here.
+[…]
+Motivating Next-Gen Accelerators with Flexible $N:M$ Activation Sparsity via Benchmarking Lightweight Post-Training Sparsification Approaches
+Shirin Alanova 1, Kristina Kazistova 1, Ekaterina Galaeva 1, Alina Kostromina 1, Vladimir Smirnov 1, Dmitry Redko 1, Alexey Dontsov 1, Maxim Zhelnin 1, Evgeny Burnaev 1,2, Egor Shvetsov 1 1 Skolkovo Institute of Science and Technology 2 Artificial Intelligence Research Institute 3 HSE University indicates equal contribution.
 1 Introduction
-The expanding capabilities of Large Language Models (LLMs) have driven enormous demand for efficient AI inference.
-A common rule of thumb states that inference serving speed for a dense model with $N$ parameters scales as $\propto 1/\sqrt{N}$ (Erdil, 2025).
-Accelerating inference typically involves either reducing numerical precision via quantization to mitigate memory bottlenecks and enable faster arithmetic (Shvetsov et al., 2024) or sparsification to reduce the number of parameters (Maximov et al., 2025).
-Sparsity improves LLM efficiency in two key ways: (1) by reducing computation, and (2) by reducing I/O traffic for transferring parameters between memory and compute units – a major bottleneck during inference.
+Large Language Models (LLMs) have intensified demand for efficient inference.
+A common rule of thumb suggests serving speed for a dense $N$-parameter model scales as $\propto 1/\sqrt{N}$ (Erdil, 2025).
+Inference is often accelerated via quantization (Frantar et al., 2022; Shvetsov et al., 2024) or sparsification (Frantar and Alistarh, 2023; Maximov et al., 2025), with sparsity reducing both compute and memory I/O.
 Weights vs.
 Activations.
-While sparse weights and activations offer identical theoretical FLOPs (floating point operations count), their practical implications differ.
-Weight sparsity enables static compression, efficiently reducing model storage and memory bandwidth.
-However, its static nature often causes irreversible model degradation.
-In contrast, activation sparsity is dynamic and input-adaptive, preserving model capacity, which we demonstrate in our work.
-Sparsity in LLMs.
-Naturally, pruning is most effective when the values to be pruned already have low magnitudes or are zero, which is often the case for some LLMs’ intermediate representations (Liu et al., 2023), especially due to the ReLU activation function in MLP blocks (Mirzadeh et al., 2023).
-Even if a model was trained with a non-ReLU activation function, we can still perform relufication Song et al.
-(2024b) to induce sparsity.
-This makes activation sparsification a natural choice for enhancing the model’s performance.
+Although weight and activation sparsity yield the same theoretical FLOP count, they differ in practice.
+Weight sparsity enables static compression but can irreversibly degrade model quality, whereas activation sparsity is input-adaptive and better preserves model capacity though it requires an additional pruning step for each input.
 Accelerating LLMs with Sparse Activations.
-Indeed, recent results demonstrated that activation sparsification supported by specialized software can increase the model’s vector by matrix multiplication during the LLM decoding stage up to $2\timestimes$ (Song et al., 2024b; a; Liu et al., 2024; Lee et al., 2024).
-The idea behind such approaches is relatively straightforward - if a vector has many zeroes, we can disregard corresponding columns in a weight matrix $W$ on the fly.
-However, as batch size grows beyond oneone, it becomes challenging to gain benefits (Shrestha et al., 2025).
-Moreover, most previous methods mainly rely on sparsity induced only by activation functions (Mirzadeh et al., 2023) in MLP blocks.
+Naturally, pruning is most effective when the values to be pruned already have low magnitudes or are zero, which is often the case for some LLMs’ intermediate representations (Liu et al., 2023; Li et al., 2022).
+Activation sparsity was employed to accelerate the decoding stage by enabling faster sparse vectordense matrix multiplications up to 2× speedup using specialized kernels (Song et al., 2024b, a; Liu et al., 2024; Lee et al., 2024).
+These gains are most pronounced for batch size 1 and diminish as batch size increases (Shrestha et al., 2025).
+The speedup primarily arises because rows of the dense matrix corresponding to zero elements in the sparse vector can be skipped during computation.
+Moreover, these methods often require some predictive mechanism to upload required weight indices into memory ahead of time Liu et al.
+(2023).
+Semi-structured N:M activation sparsity, keeping $N$ non-zeros per block of size $M$, can extend benefits beyond single vector–matrix products by employing hardware support.
+Although 2:4 activation sparsity has been explored for training (Haziza et al., 2025; Wang et al., 2024), and post-training weight sparsification Maximov et al.
+(2025); Han et al.
+(2015); Frantar and Alistarh (2023); Kurtić et al.
+(2023) post-training N:M activation pruning is largely unexplored because it is not supported on commercial hardware.
 […]
 Focus and Motivation.
-Applying N:M sparsity patterns to activations, while conceptually straightforward, raises several key questions: (Q1) Selection Strategy: As with weight sparsification, the criterion for selecting which activations to prune is crucial for maintaining performance (Zhelnin et al., 2024).
-(Q2) Error Mitigation: Fine-tuning is a standard method for recovering performance lost to post-training sparsification, but can be prohibitively expensive or risk compromising model safety alignment (Kharinaev et al., 2025).
-Therefore, we focus on lightweight, plug-and-play and hardware-friendly approaches that require only minimal calibration data (e.g., WikiText) or no data at all to enhance performance.
-(Q3) Beyond 2:4 Patterns: While 2:4 is the only hardware-supported semi-structured pattern, we explore a broader range {2:4,4:8,8:16,16:32} to motivate future hardware design.
-While we cannot directly compare acceleration gains between various patterns, our work solely focuses on model performance.
-Our Contributions.
-This paper makes the following key contributions:
-  • Activation vs.
-Weight Sparsity Superiority: We show that, at similar sparsity levels, activation pruning consistently outperforms weight pruning across four diverse LLMs: Llama2-7B, Llama3-8.1B, Qwen2.5-7B, and Gemma3-4B.
-  • Lightweight Error Mitigation and Selection Criteria: We comprehensively evaluate four plug-and-play error mitigation techniques, three evaluated for the first time in the context of activation sparsity.
-We also evaluate three pruning selection criteria.
-These methods include statistical criteria such as median shift and variance correction and provide strong, hardware-friendly baselines for activation sparsification.
-[…]
-2 Hardware and Comparison of 2:4 and 8:16 Sparsity
-One of the primary aims of our work is to motivate the design of hardware for more flexible sparsity patterns.
-Currently, only NVIDIA GPUs support structured sparsity, specifically the 2:4 pattern.
-This pattern achieves  1.5–1.7x inference acceleration and  1.5x energy reduction for 7B models, primarily from a 2x reduction in memory bandwidth (Fang et al., 2024; Lin et al., 2023).
-At the same time, any N:M pattern beyond 2:4, including 8:16, can provide the same 2x bandwidth reduction but with greater flexibility.
-In the general N:MN{:}M format, 2:4 sparsity permits $\binom{4}{2}=6$ non-zero patterns per block, requiring $\log_{2}(6)/4\approx 0.75$ bits per element for metadata.
-In contrast, the 8:16 pattern allows $\binom{16}{8}=12{,}870$ distinct configurations, demanding $\log_{2}(12{,}870)/16\approx 0.875$ bits per element.
-Stacking four 2:4 blocks to form a 16-element unit yields only $6^{4}=1{,}296$ configurations—nearly an order of magnitude fewer than the 12,870 offered by a native 8:16 block.
-Although 8:16 requires slightly more metadata storage, its vastly greater configurational flexibility makes it a compelling candidate.
-The main trade-off is implementation complexity.
-Decoding sparsity metadata and gathering non-zero values requires sophisticated circuitry.
-As the block size $M$ grows, these gather operations could impact performance.
-However, operating on larger, aligned blocks could also improve cache utilization, potentially amortizing this overhead.
+The primary goal of this work is twofold: to motivate the development of hardware that natively supports semi-structured activation sparsity, and to benchmark existing approaches for inducing such sparsity in activations.
+Generally, these approaches address two key challenges:
+(P1) Selection Strategy. As with weight sparsification, the choice of which activations to retain critically affects model accuracy (Zhelnin et al., 2025).
+(P2) Error Mitigation.
+While post-training fine-tuning can recover lost performance, it is often impractical due to computational cost or risks to safety alignment (Kharinaev et al., 2025).
+We therefore mainly focus on lightweight, plug-and-play methods that require minimal (e.g., WikiText) or zero calibration data.
+Our Contributions:
+  • First, we demonstrate that activation sparsity outperforms weight sparsity.
+At matched sparsity levels, activation pruning consistently yields higher accuracy than weight pruning across four diverse large language models Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct highlighting activations as a more promising target for future sparse accelerators.
+  • Second, we benchmark four plug-and-play error mitigation techniques, three of which are applied to semi-structured activation sparsity for the first time, including statistical approaches such as median shift and variance correction.
+We also evaluate three selection criteria.
+These methods establish strong, retraining-free baselines that require minimal metadata, aligning well with hardware constraints.
+  • Finally, we analyze a wide range of structured sparsity patterns and show that larger patterns dramatically improve model fidelity: the 16:32 pattern achieves performance close to unstructured 50% sparsity and retains over three times more accuracy than the conventional 2:4 pattern.
+Nevertheless, considering implementation trade-offs, we advocate for 8:16 as the optimal balance it offers twice the accuracy retention of 2:4 while remaining highly practical for near-term hardware adoption.
+Together, these results provide concrete evidence that expanding hardware support beyond static 2:4 weight sparsity can unlock significant efficiency gains without compromising model quality, thereby motivating the next generation of sparsity-aware architectures.
+2 Methods
+In the sections below, we formally define the pruning criteria and error mitigation strategies evaluated in this work.
+A brief summary of these methods is provided in Table 1 and more detailed Table is available in Appendix 6.
+Table 1: Summary of evaluated pruning metrics and transformations. Full descriptions are provided in Appendix 6.
+Short name Description
+Pruning metrics
+ACT Activation magnitude
+WT Weight magnitude
+CLACT* Context-aware cosine-loss-inspired
+Amber-Pruner Weight-informed importance
+Transformations
+D-PTS* Dynamic per-token shift
+S-PTS Static per-token shift
+L-PTS* Learnable per-token shift
+VAR* Variance correction
+VAR+L-PTS* VAR + learnable shift
+R-Sparse Sparse activations + low-rank weights
+2.1 Preliminaries
+For a linear layer with weights $\mathbf{W}$, input activations $\mathbf{X}$, and output $\mathbf{Y}$:
+$\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}.$ (1)
+We construct a sparsity mask $\mathbf{M}$ using a metric $S$ and threshold $t$:
+$\mathbf{M}_{ij}=\begin{cases}1,&S(\mathbf{X}_{ij})\geq t,\\ 0,&S(\mathbf{X}_{ij})<t,\end{cases}$ (2)
+yielding
+$\mathbf{Y}_{p}=(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}.$ (3)
+Unstructured sparsity applies a global threshold over all elements.
+Semi-structured N:MN{:}M sparsity partitions rows (or columns) into non-overlapping blocks of size $M$ and keeps the top-$N$ elements per block by $S$ (e.g., 2:4 removes 50%) (Hu et al., 2024).
+We also study 8:16, which retains the same 50% density with higher flexibility and modest metadata overhead (Maximov et al., 2025).
+2.2 Pruning Criterion
+ACT: This is the magnitude activation pruning metric (ACT), defined as the absolute value of the element $\mathbf{X}_{ij}:S_{\textit{ACT}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|$.
+WT: This is the weight-based pruning metric (WT), defined as the absolute value of the corresponding weight $\mathbf{W}_{ij}$: $S_{\textit{WT}}(\mathbf{W}_{ij})=|\mathbf{W}_{ij}|$.
+CLACT: Inspired by output cosine similarity in (Mi et al., 2025), we propose Cosine Loss ACTivation (CLACT), a context-aware score that emphasizes activations aligned with their row/column energy:
+$S_{\textit{CLACT}}(\mathbf{X}_{ij})=\dfrac{|\mathbf{X}_{ij}|}{\sqrt{\sum_{k=1}^{h}\mathbf{X}^{2}_{ik}}}\,\sqrt{\sum_{p=1}^{l}\mathbf{X}^{2}_{pj}},$ (4)
+where $h$ is the hidden dimension and $l$ is the sequence length.
+Amber-Pruner: Following (An et al., 2025), we (i) remove weight outliers outside the 0.5–99.5 percentiles, (ii) standardize the remaining weights, and (iii) score each activation as $S_{\textit{Amb.-Pr.}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|\cdot\mathcal{L}(\hat{\mathbf{W}}_{:,j})$, where $\mathcal{L}(\cdot)$ is the channel-wise $\ell_{2}$ norm.
+CLACT is context-aware (and for l=1l{=}1 reduces to an $\ell_{1}$-type criterion), whereas Amber-Pruner leverages weight magnitudes but is not context-aware.
+2.3 Transformations for Error Mitigation
+D-/S-/L-PTS (dynamic/ static/ learnable per-token shift) centers activations near zero: $\hat{\mathbf{X}}=\mathbf{X}-\bm{\eta}$, and uses the compensated form $\mathbf{Y}_{p}=((\hat{\mathbf{X}}\odot\mathbf{M})+\bm{\eta})\mathbf{W}^{\top}$ (Chua et al., 2024), where $\bm{\eta}=\overline{\mathbf{X}}$ for D-PTS; S-PTS uses a fixed $\bm{\eta}$ collected in a short warm-up; L-PTS learns $\bm{\eta}$.
+VAR applies per-token variance correction after pruning: $\mathbf{Y}_{p}=\bm{\nu}(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}$, with
+$\nu=\sqrt{\dfrac{\operatorname{Var}[\mathbf{X}]}{\operatorname{Var}[\mathbf{X}\odot\mathbf{M}]}}.$ (5)
+VAR+L-PTS combines VAR scaling with the learnable shift. R-Sparse combines activation sparsity with a low-rank weight approximation (Kamirul et al., 2025) (details in Appendix B).
+For methods requiring calibration/learning, we use WikiText-2: S-PTS stores a fixed bias vector $\bm{\eta}$, L-PTS (and VAR+L-PTS) learns $\bm{\eta}$, and R-Sparse learns low-rank factors; all other methods require no learned parameters.
+2.4 Evaluation & Models
+We evaluate the methods in Table 1 in two stages.
+First, we use Core Datasets (BoolQ, WinoGrande, PIQA, ARC-Easy) to screen all methods.
+We then focus on the most promising approaches on Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct, and extend evaluation to Extended Datasets (HellaSwag, OpenBookQA, RTE, MMLU, Lambada_standard, Lambada_openai, IFEval).
+When calibration is required, we use WikiText-2, following common compression practice Egiazarian et al.
+(2024); Frantar et al.
+(2022); Van Baalen et al.
+(2024).
+All results are obtained with LM Eval Harness Gao et al.
+(2023); dataset details are in Table 9.
+We report results for Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct.
+For Qwen2.5-7B-Instruct, we do not sparsify key/query/value activations due to severe degradation observed in preliminary experiments.
 Figure 1: Comparison of unstructured sparsity in activations (ACT) and weights (WT) averaged across four datasets at varying sparsity ratios.
@@
 Figure 2: Comparison of sparsity patterns with unstructured sparsity. 50% and 70% correspond to unstructured sparsity. Lower is Better. More detailed results are presented in Appendix Table 7.
-3 Methods
-Table 1: Briefly describe evaluated activation pruning metrics (top) and transformations (bottom).
+3 Results
+3.1 Sparse Weights vs. Activations
+In Figure 1 and Table 10, we demonstrate that unstructured weight sparsification causes greater model degradation than unstructured activation sparsification at the same sparsity levels: {20%, 50%, 70%, 90%}.
+For this evaluation, we specifically use unstructured magnitude-based sparsification, as it is less damaging than semi-structured sparsification and thus provides a lower bound on performance degradation.
+3.2 Optimal semi-structured sparsity patterns
+Our preliminary investigation demonstrates that while the 16:32 pattern achieves performance close to unstructured sparsity (a 5.4% drop versus 4.5% for 50% unstructured), it requires more metadata and greater resources for gather operations, as discussed in Section 1.
+Therefore, we focus on the 8:16 pattern, despite its higher performance drop of 7.38%.
+For comparison, the 2:4 pattern results in a 14.35% drop.
+These results are shown in Figure 2 and Table 7 in Appendix, we used only magnitude pruning to obtain these results.
+By demonstrating the superior model quality of 8:16 sparsity, our work incentivizes hardware designers to invest in or at least consider supporting the 8:16 pattern.
+Table 2: Average relative performance drop (%, lower is better; negative = improvement) across four datasets, averaged over Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct.
+Act/Wt: activation/weight sparsity.
+Selection: ACT/CLACT/Amber-Pruner; transformations: VAR, D-PTS, S-PTS, L-PTS, R-SPARSE.
+Methods marked with an asterisk (*) are proposed in this paper.
+Full results: Appendix 11 and  12.
+Target Pattern Method Avg drop ($\downarrow$)
+Act u50 ACT $3.82$
+Wt 2:4 WT $24.49$
+Act 2:4 ACT $9.67$
+Act 2:4 CLACT* $7.79$
+Act 2:4 Amber-Pruner $7.85$
+Act 2:4 VAR* $6.09$
+Act 2:4 D-PTS* $5.84$
+Act 2:4 S-PTS $4.29$
+Act 2:4 L-PTS* $8.79$
+Act 2:4 R-SPARSE (64) $7.70$
+Act 2:4 R-SPARSE (128) $8.05$
+Wt 8:16 WT $17.68$
+Act 8:16 ACT $5.47$
+Act 8:16 CLACT* $2.29$
+Act 8:16 Amber-Pruner $1.56$
+Act 8:16 VAR* $3.30$
+Act 8:16 D-PTS* $2.07$
+Act 8:16 S-PTS $0.61$
+Act 8:16 L-PTS* $5.32$
+Act 8:16 R-SPARSE (64) $1.52$
+Act 8:16 R-SPARSE (128) $2.63$
+3.3 Results on Single/Multi-choice Datasets
+3.3.1 Evaluation of Pruning Criteria
+We evaluate CLACT, Amber-Pruner, and magnitude pruning as a baseline.
+The main results for the 2:4 and 8:16 sparsity patterns are presented in Table 2.
+On average, both CLACT and Amber-Pruner outperform magnitude pruning by at least 2%, however, we observe no clear winner between them.
+As noted in Section 2.2, these criteria are designed for different purposes: CLACT adjusts based on context, while Amber-Pruner adjusts based on weight magnitudes.
+Notably, for Llama3.1-8B-Instruct under the 2:4 sparsity pattern, simple magnitude pruning outperforms both advanced criteria, underscoring model and architecture-specific sensitivities to pruning strategies.
+3.3.2 Evaluation of Transformations
+Our main results are presented in Table 2.
+Surprisingly, we find that simple methods such as dynamic and static per-token shifts (D-PTS, S-PTS) outperform most other approaches.
+The second most effective methods are VAR and R-SPARSE.
+We also observe that increasing the number of dimensions in R-SPARSE (from 64 to 128) leads to worse performance, which may indicate overfitting on the calibration data.
+Finally, we note that L-PTS, the approach with learnable per-token shifts, significantly underperforms compared to its static counterpart, S-PTS.
+3.4 Instruction-Following Tasks
+Table 3 presents instruction-following performance on the IFEval benchmark for Llama3.1-8B-Instruct and Qwen2.5-7B-Instruct, evaluated under two semi-structured sparsity patterns (2:4 and 8:16) and four activation transformation methods: S-PTS, D-PTS, R-Sparse, and VAR.
+First of all, we observe a strong model degradation on generative tasks.
+Second of all, we see that VAR is the strongest performer overall, especially for Llama3.1-8B-Instruct.
+S-PTS/D-PTS are competitive and lightweight, and R-Sparse lags significantly, particularly at 2:4.
+We speculate that while semi-structured patterns are good for prefill stage in LLMs they significantly degrade performance during decode stage.
+However, as discussed in Section 1 decode stage for single vector can be accelerated with more flexible approaches.
+Table 3: Instruction-following (IFEval) prompt-level accuracy reported as PS/PL (prompt-level strict/loose accuracy).
+Model Method 2:4 8:16
+Llama3.1-8B ORIG 0.4455/0.4861 0.4455/0.4861
+S-PTS 0.1682/0.1904 0.2995/0.3327
+D-PTS 0.1941/0.2015 0.2828/0.3198
+R-Sparse 0.0869/0.0979 0.2089/0.2311
+VAR 0.2237/0.2458 0.3161/0.3586
+Qwen2.5-7B ORIG 0.7135/0.7394 0.7135/0.7394
+S-PTS 0.4325/0.5176 0.5194/0.5804
+D-PTS 0.4399/0.5120 0.5434/0.5989
+R-Sparse 0.2736/0.3457 0.3697/0.4196
+VAR 0.4565/0.5342 0.5249/0.5896
+3.5 Results for Unstructured Pruning
+We evaluate D-PTS, VAR, and two selection criteria in this experiment: CLACT and Amber-Pruner using the Llama3.1-8B-Instruct model.
+Results in Table 4 indicate that VAR is the most effective transformation under unstructured sparsity.
+Moreover, CLACT outperforms Amber-Pruner by a wider margin here than in our semi-structured pruning experiments.
+These findings suggest two key insights: (1) No single method emerges as optimal for both unstructured and semi-structured sparsity.
+(2) The methods proposed in this work, VAR and CLACT, demonstrate strong generalization and are well-suited for both semi-structured and unstructured activation pruning.
+Table 4: Performance comparison of pruning methods under 50% and 70% unstructured sparsity for Llama3.1-8B-Instruct model.
+Method ArcE BoolQ PIQA Wino Grande (%) Avg. Drop
+Original $0.821$ $0.839$ $0.800$ $0.734$
+Unstructured 50%
+ACT $0.777$ $0.820$ $0.771$ $0.686$ $4.450$
+D-PTS $0.786$ $0.825$ $0.781$ $0.690$ $3.600$
+VAR $0.784$ $0.819$ $0.776$ $0.705$ 3.470
+CLACT $0.780$ $0.825$ $0.766$ $0.700$ $3.890$
+Amber $0.768$ $0.820$ $0.763$ $0.702$ $4.450$
+Unstructured 70%
+ACT $0.558$ $0.631$ $0.647$ $0.548$ $25.320$
+D-PTS $0.565$ $0.624$ $0.651$ $0.534$ $25.680$
+VAR $0.614$ $0.651$ $0.676$ $0.532$ 22.660
+CLACT $0.555$ $0.604$ $0.627$ $0.524$ $27.670$
+Amber $0.487$ $0.594$ $0.590$ $0.539$ $30.680$
+3.6 Combination of Methods
+Next, we evaluated combinations of multiple approaches to explore potential performance gains.
+These combinations and their results are presented in Table 8.
+As shown, none of the evaluated combinations outperforms any single method, highlighting the challenges of naively combining them.
+Table 5: Llama3.1-8B-Instruct with 8:16 activation sparsity: aggregated results across layer subsets. ORIG. AVG. = 0.6726. Drop is computed w/o perplexity. Full results: Appendix 13.
+Method Layers PPL AVG. Drop $\downarrow$
+LS+L-PTS all 9.6036 0.6047 10.90%
+LS+L-PTS key,out,gate,down 8.3483 0.6385 5.43%
+LS+L-PTS key,value,gate,down 8.0821 0.6503 3.56%
+LS+L-PTS+VAR all 9.4983 0.6056 10.60%
+LS+L-PTS+VAR key,out,gate,down 8.2930 0.6422 4.64%
+LS+L-PTS+VAR key,value,gate,down 8.0259 0.6516 3.36%
+3.7 Layer Sensitivity
+We study layer sensitivity to 8:16 activation sparsity on Llama3.1-8B-Instruct using the Extended Datasets (Table 5).
+This analysis focuses on learnable methods, mainly L-PTS and LS (learnable diagonal scaling).
+Although learnable approaches underperform on average in our broader experiments, the best 8:16 configurations for Llama3.1-8B-Instruct rely on learnable parameters (Table 5).
+We observe that the FFN up projection and the attention out projection [^1] The output projection of the attention mechanism.
+It combines outputs from all attention heads and projects them back to the model’s hidden dimension.
+are the most sensitive: pruning them causes the largest drops, suggesting they should be preserved or handled with extra care.
+While this may not generalize to all layers, it indicates that layer importance under activation sparsity is highly non-uniform.
+4 Discussion
+The performance gap between multiple/single-choice benchmarks (e.g., BoolQ, PIQA) and IFEval likely stems from differences in the inference stages they emphasize.
+Core QA benchmarks primarily stress the prefill phase, whereas IFEval evaluates both prefill and autoregressive generation.
+Our evaluation remains valid: semi-structured patterns like 2:4 and 8:16 are especially effective at accelerating the prefill stage, which often dominates inference latency.
+5 Conclusion
+This work establishes that post-training activation pruning is significantly more accuracy-preserving than weight pruning in large language models.
+Across four diverse architectures (Llama2-7B, Llama3.1-8B, Qwen2.5-7B, and Gemma3-4B), we demonstrate that activation sparsity consistently retains model capabilities better than weight sparsity at matched sparsity levels.
+Our evaluation reveals that lightweight error mitigation techniques particularly CLACT, D-PTS, and VAR establish strong, hardware friendly baselines requiring minimal calibration data.
+Through systematic analysis of semi-structured patterns, we find that 16:32 approaches unstructured 50% sparsity in fidelity, while 8:16 emerges as the optimal near term target.
+6 Limitations
+Key limitations: First, all evaluations use software emulation without hardware measurements of speedup or energy efficiency.
+Second, layer sensitivity analysis remains preliminary; while FFN up-projections and attention out-projections appear most vulnerable, broader architectural studies are needed.
+Third, generative performance (IFEval) degrades significantly more than multiple-choice QA under sparsity, revealing an evaluation bias toward prefill-dominated workloads.
+The anomalous improvements on Qwen2.5-7B-Instruct for some benchmarks further highlight dataset-specific artifacts rather than genuine capability preservation.
+Our hardware overhead analysis in Appendix A is a rough estimate and more precise analysis is required.
+Appendix
+Appendix A Hardware Implications and Computational Overhead Analysis
+While our empirical results demonstrate significant accuracy benefits from flexible N:M activation sparsity, the practical value of these techniques depends critically on whether hardware implementations can overcome the computational overhead of dynamic sparsification.
+This section provides a comprehensive break-even analysis to determine the hardware conditions required for activation sparsity to deliver net performance and efficiency gains.
+The fundamental challenge for activation sparsity is that the theoretical memory bandwidth reduction must overcome the overhead introduced by dynamic sparsification operations.
+Based on literature and performance models, we estimate break-even thresholds where benefits begin to outweigh costs.
+A.1 Energy-Delay Product Analysis
+The Energy-Delay Product (EDP) provides a comprehensive metric for evaluating whether activation sparsity delivers net efficiency benefits.
+For 8:16 sparsity to be worthwhile from an EDP perspective, it must overcome both computational overhead and energy costs of the sparsification process itself.
+We model EDP improvement for semi-structured sparsity patterns as:
+$\text{EDP}_{\text{improvement}}=\frac{\text{EDP}_{\text{dense}}}{\text{EDP}_{\text{sparse}}}\approx\frac{r\cdot\eta}{1+\alpha}$
+Where:
+  • $r=2.0$ is the theoretical bandwidth reduction ratio for 8:16 sparsity
+  • $\eta=0.85$ is the hardware utilization efficiency (representative of practical implementations)
+  • $\alpha=0.3$ is the overhead factor from sparsification operations
+This overhead factor $\alpha$ is calibrated from real measurements Fang et al.
+(2024) demonstrated that dynamic activation sparsification incurs 30–35% additional latency on current hardware without native support, broken down as:
+  • Activation magnitude computation and block-wise selection
+  • Mask application and metadata handling
+  • Error mitigation techniques (D-PTS, VAR), which is not included in Fang et al. (2024).
+Solving for the minimum hardware acceleration factor $k$ required for net EDP benefits:
+$r\cdot\eta>k\cdot(1+\alpha)$
+$2.0\cdot 0.85>k\cdot(1+0.3)$
+$k>\frac{1.7}{1.3}\approx 1.31$
+However, due to our imprecise estimations we will consider a higher amortized $k>1.6\times$ required for speedup.
+A.2 Hardware Implementation Requirements
+[…]
+  • Dedicated sparsity controllers: On-chip circuitry that can generate sparsity masks with minimal latency, reducing the 15–20% selection overhead
+  • Hardware-supported statistical units: Specialized units for variance/mean calculation required by error mitigation techniques, eliminating their computational overhead
+  • Hierarchical sparsity support: Different patterns for different layer types based on sensitivity analysis (Section 3.7)
+  • Bandwidth-optimized gather operations: Specialized memory controllers that maintain high efficiency despite irregular access patterns
+Recent hardware developments show promising directions: NVIDIA’s Blackwell architecture includes a dedicated hardware decompression engine Jarmusch and Chandrasekaran (2025), while the LazyGPU microarchitecture enables lazy memory request issuing Liu et al.
+(2025).
+[…]
+Table 6: Brief description of evaluated activation pruning metrics (top) and transformations (bottom).
 Abbreviations in bold with an asterisk (*) denote methods proposed here or first evaluated with sparse activations.
@@
 CLACT* Cosine Loss Activation A metric inspired by cosine similarity from Mi et al. (2025)
-Amber-Pruner (2025a) Weights Importance A metric which accounts for important weights after outlier removal and normalization
+Amber-Pruner (2025) Weights Importance A metric which accounts for important weights after outlier removal and normalization
 Transformations
@@
 R-Sparse (2025) Rank-Aware Sparsity Combines sparse activations with weight low-rank SVD factors learned on WikiText-2
-3.1 Preliminaries
-Consider a linear layer of a LLM with weight $\mathbf{W}$, input activations $\mathbf{X}$ and output $\mathbf{Y}$ then:
-$\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}$ (1)
-Our aim is to find a semi-structured pruning mask $\mathbf{M}$ for the activations $\mathbf{X}$ based on some pruning metric $S$, such that:
-[…]
-where $i,j$ denote the indices of the matrix elements, $\mathrm{t}$ is a threshold value chosen to ensure the desired level of sparsity.
-Once the mask $\mathbf{M}$ is computed, the output of the linear layer with pruned activations is given by:
-$\mathbf{Y}_{p}=(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}$ (3)
-[…]
-3.2 Pruning Criterion
-In this study, we primarily consider some pruning metrics. You can find brief descriptions of them in 1.
-ACT: This is the magnitude activation pruning metric (ACT), defined as the absolute value of the element $\mathbf{X}_{ij}:S_{\textit{ACT}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|$.
-WT: This is the weight-based pruning metric (WT), defined as the absolute value of the corresponding weight $\mathbf{W}_{ij}$: $S_{\textit{WT}}(\mathbf{W}_{ij})=|\mathbf{W}_{ij}|$.
-CLACT: This metric is inspired by (Mi et al., 2025), which measures the cosine similarity between layer outputs.
-Building on this concept, we propose a novel metric called Cosine Loss ACTivation (CLACT).
-This approach approximates the cosine distance between an activation $\mathbf{X}_{ij}$ and its feature vector, quantifying each element’s importance within its activation row.
-Unlike magnitude-based methods, CLACT identifies activations that are contextually significant rather than merely large in absolute value.
-$S_{\textit{CLACT}}(\mathbf{X}_{ij})=\dfrac{|\mathbf{X}_{ij}|}{\sqrt{\sum_{k=1}^{h}\mathbf{X}^{2}_{ik}}}\sqrt{\sum_{p=1}^{l}\mathbf{X}^{2}_{pj}},$ (4)
-where $h$ is hidden dimension, $l$ is sequence length, $j$ is a token number and $i$ an element withing this token.
-Amber-Pruner: This metric from An et al.
-(2025b) consists of the following steps: (1) Outlier Removal.
-To reduce the influence of extreme values, weights $\mathbf{W}$ outside the 0.5th99.5th0.5th-99.5th percentiles are discarded: $\mathbf{W}=\left\{\omega_{k}\mid Q_{0.005}(\mathbf{W})\leq\omega_{k}\leq Q_{0.995}(\mathbf{W})\right\}$.
-(2) Normalization.
-The remaining weights are standardized using their mean and variance: $\hat{\mathbf{W}}_{ij}=(\mathbf{W}_{ij}-\mathbb{E}[\mathbf{W}])(\operatorname{Var}[\mathbf{W}])^{-\frac{1}{2}}$.
-(3) Channel-wise Scoring.
-For each channel $\hat{\mathbf{W}}_{:,j}$, the $\ell_{2}$ norm is computed, and final scores are assigned to activation elements, $\mathcal{L}(\cdot)$ - denotes channel-wise $\ell_{2}$ norm: $S_{\textit{Amb.-Pr.}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|\cdot\mathcal{L}\left(\hat{\mathbf{W}}_{:,j}\right)$.
-Here, we highlight a critical distinction between CLACT and Amber-Pruner.
-CLACT is a context aware criterion, particularly useful during the prefill stage since with $l=1$, it converges to the $L1$ - norm criterion, moreover, it does not consider weight magnitudes.
-In contrast, Amber-Pruner explicitly prioritizes weights based on their values and remains effective across both prefill and generation stages but does not consider context.
-3.3 Error mitigation and Transformations
-Since semi-structured pruning of activations can significantly disrupt the outputs of linear layers, additional transformations can be applied to mitigate these negative effects and restore the pruned activations to their original distribution.
-You can find brief descriptions of them in 1.
-PCS: This approach is inspired by SmoothQuant (Xiao et al., 2023) and has been used for 2:4 weight sparsification (Maximov et al., 2025) as well.
-It uses a per-channel smoothing (PCS) factor $s$ applied to activations $\mathbf{X}$ before pruning: $\hat{\mathbf{X}}=\mathrm{diag}(s(\mathbf{X},\mathbf{W}))^{-1}\mathbf{X}$.
-To compensate for this transformation, the weight matrix is correspondingly scaled in the opposite direction.
-As a result, the output of the linear layer is computed as: $\mathbf{Y}_{p}=(\hat{\mathbf{X}}\odot\mathbf{M})(\mathrm{diag}(s(\mathbf{X},\mathbf{W}))\mathbf{W}^{\top})$.
-The scale factor $s$ is estimated as Xiao et al.
-(2023): $s=\sqrt{\max|\mathbf{X}_{:,j}|(\max|\mathbf{W}^{\top}_{:,j}|)^{-1}}$
-D-/S-/L-PTS: The next approach applies a dynamic per-token shift (D-PTS) to the input activations $\mathbf{X}$ using a bias term $\mathbf{\eta}$, such that $\hat{\mathbf{X}}=\mathbf{X}-\mathbf{\eta}$.
-The purpose of this shift is to center most elements of the activations near zero.
-As a compensatory transformation, the bias $\eta$ is added to the weights, so the output of the linear layer is given by Chua et al.
-(2024): $\mathbf{Y}_{p}=((\hat{\mathbf{X}}\odot\mathbf{M})+\mathbf{\eta})\mathbf{W}^{\top}$, where $\eta=\overline{\mathbf{X}}$.
-A related variant is the static per-token shift (S-PTS), where the bias $\eta$ is collected during a short warm-up phase and then fixed for all subsequent batches.
-This reduces computational overhead while providing similar centering effects.
-Finally, we introduce a learnable per-token shift (L-PTS).
-VAR: In contrast to the previous approach, this method applies a per-token scaling factor $\mathbf{\nu}$ to pruned activations such that the output matrix is given as: $\mathbf{Y}_{p}=\mathbf{\nu}(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}$, The scale factor $\mathbf{\nu}$ is determined as
-$\nu=\sqrt{\dfrac{\operatorname{Var}[\mathbf{X}]}{\operatorname{Var}[\mathbf{X}\odot\mathbf{M}]}}$ (5)
-We refer to this method as variance correction VAR.
-VAR+L-PTS: This method is a combination of the per-token scaling with factor VAR $\mathbf{\nu}$ is given by Eq. 5 and the shift L-PTS.
-R-Sparse: This method from  (Kamirul et al., 2025) combines activation sparsity with a low-rank approximation of the weight matrix. We provide more details in Appendix A.
-Throughout our experiments we use the WikiText-2 dataset for calibration and learning some parameters.
-For S-PTS we calibrate a single per-channel bias vector $\eta$ and then keep it fixed for all subsequent batches.
-For the learnable variant L-PTS (and the combined VAR+L-PTS) we train $\eta$ and fine-tune other weights.
-For R-Sparse we train in the same way the low-rank factors of the weight adaptation.
-In all other methods no parameters are learned or calibrated.
-3.4 Evaluation & Models
-We begin with a broad funnel of experiments to evaluate all methods listed in Table 1.
-At this initial stage, we use the Core Datasets, Single-Choice and Multi-Choice benchmarks including BoolQ, WinoGrande, PIQA, and ARC-Easy.
-Once this exploratory phase is complete, we narrow our focus to the most promising approaches, evaluating them primarily on two models: Qwen2.5-7B-Instruct and Llama3-8B-Instruct.
-This more targeted evaluation phase not only refines the set of methods under consideration but also expands the evaluation to the Extended Datasets: HellaSwag, OpenBookQA, RTE, MMLU, Lambada_standard, Lambada_openai, and IFEval.
-For calibration tasks, when required, we use WikiText-2.
-This dataset selection follows established model compression practices Egiazarian et al.
-(2024); Frantar et al.
-(2022); van Baalen et al.
-(2024).
-All evaluations are performed using the LM Eval Harness Gao et al.
-(2023), and full dataset details are provided in Table 8.
-For the models we focus on: Qwen2.5-7B-Instruct, Llama3-8B-Instruct, Llama2-7B, Gemma3-4B-Instruct.
-Importantly, for Qwen2.5-7B-Instruct, we did not sparsify the key, query, or value activations, as preliminary experiments showed severe model degradation when these layers were pruned.
-4 Results
-4.1 Sparse Weights vs. Sparse Activations
-In Figure 2 and Table 9, we demonstrate that unstructured weight sparsification causes greater model degradation than unstructured activation sparsification at the same sparsity levels: {20%, 50%, 70%, 90%}.
-For this evaluation, we specifically use unstructured magnitude-based sparsification, as it is less damaging than semi-structured sparsification and thus provides a lower bound on performance degradation.
-4.2 Optimal semi-structured sparsity patterns
-Our preliminary investigation demonstrates that while the 16:32 pattern achieves performance close to unstructured sparsity (a 5.4% drop versus 4.5% for 50% unstructured), it requires more metadata and greater resources for gather operations, as discussed in Section 2.
-Therefore, we focus on the 8:16 pattern, despite its higher performance drop of 7.38%.
-For comparison, the 2:4 pattern results in a 14.35% drop.
-These results are shown in Figure 2 and Table 6 in Appendix, we used only magnitude pruning to obtain these results.
-By demonstrating the superior model quality of 8:16 sparsity, our work incentivizes hardware designers to invest in—or at least consider—supporting the 8:16 pattern.
-Table 2: Average relative performance (%) across four datasets.
-Values indicate performance drops (lower is better), negative values signify performance improvement.
-Methods marked with an asterisk (*) are proposed in this paper.
-Full, non-aggregated results are available in Appendix 10.
-Models
-Method Llama2-7B Qwen2.5-7B Gemma3-4B LLama3-8B Average Drop ($\downarrow)$
-Activations Unstructured 50%
-ACT 2.312.31% 3.873.87% 4.804.80% 4.304.30% 3.823.82%
-2:4 Weight Sparsity
-WT 16.5216.52% 12.9612.96% 34.8634.86% 33.6333.63% 24.4924.49%
-2:4 Activations Selection Criteria
-ACT 9.439.43% 4.954.95% 9.949.94% 14.3514.35% 9.679.67%
-CLACT* 8.32% 2.45-2.45% 8.018.01% 17.2717.27% 7.79%
-Amber-Pruner 11.7011.70% 1.23-1.23% 5.915.91% 15.0115.01% 7.857.85%
-2:4 Activations Transformations
-VAR* 9.769.76% 1.48-1.48% 2.96% 13.1113.11% 6.096.09%
-D-PTS* 10.6710.67% -6.46% 4.584.58% 14.5914.59% 5.845.84%
-S-PTS 10.3710.37% 4.43-4.43% 3.933.93% 7.31% 4.29%
-L-PTS* 13.1313.13% 3.663.66% 4.214.21% 14.1314.13% 8.798.79%
-R-SPARSE (64) 12.9012.90% 2.55-2.55% 5.175.17% 15.2815.28% 7.707.70%
-R-SPARSE (128) 12.2312.23% 1.51-1.51% 5.165.16% 16.3416.34% 8.058.05%
-8:16 Weight Sparsity
-WT 7.847.84% 9.549.54% 26.1126.11% 27.2627.26% 17.6817.68%
-8:16 Activations Selection Criteria
-ACT 5.375.37% 4.384.38% 4.764.76% 7.387.38% 5.475.47%
-CLACT* 3.983.98% 4.02-4.02% 0.600.60% 8.608.60% 2.292.29%
-Amber-Pruner 5.325.32% 6.28-6.28% 0.080.08% 7.137.13% 1.56%
-8:16 Activations Transformations
-VAR* 4.854.85% 1.931.93% -1.87% 8.308.30% 3.303.30%
-D-PTS* 4.634.63% -8.28% 5.165.16% 6.79% 2.072.07%
-S-PTS 3.87% 7.24-7.24% 1.54-1.54% 7.37.3% 0.61%
-L-PTS* 8.158.15% 1.711.71% 4.214.21% 7.197.19% 5.325.32%
-R-SPARSE (64) 5.915.91% 6.91-6.91% 1.36-1.36% 8.448.44% 1.521.52%
-R-SPARSE (128) 7.937.93% 5.44-5.44% 0.44-0.44% 8.448.44% 2.632.63%
-4.3 Evaluation of Pruning Selection Criteria on Single/Multi-choice Datasets
-We evaluate CLACT, Amber-Pruner, and magnitude pruning as a baseline.
-The main results for the 2:4 and 8:16 sparsity patterns are presented in Table 2.
-On average, both CLACT and Amber-Pruner outperform magnitude pruning by at least 2%, however, we observe no clear winner between them.
-As noted in Section 3.2, these criteria are designed for different purposes: CLACT adjusts based on context, while Amber-Pruner adjusts based on weight magnitudes.
-Notably, for Llama3-8B under the 2:4 sparsity pattern, simple magnitude pruning outperforms both advanced criteria, underscoring model and architecture-specific sensitivities to pruning strategies.
-4.4 Evaluation of Transformations on Single/Multi-choice Datasets
-Our main results are presented in Table 2.
-Surprisingly, we find that simple methods such as dynamic and static per-token shifts (D-PTS, S-PTS) outperform most other approaches.
-The second most effective methods are VAR and R-SPARSE.
-We also observe that increasing the number of dimensions in R-SPARSE (from 64 to 128) leads to worse performance, which may indicate overfitting on the calibration data.
-Finally, we note that L-PTS, the approach with learnable per-token shifts, significantly underperforms compared to its static counterpart, S-PTS.
-4.5 Evaluation transformations with Instruction-Following Tasks
-Table 3 presents instruction-following performance on the IFEval benchmark for Llama3-8B and Qwen2.5-7B, evaluated under two semi-structured sparsity patterns (2:4 and 8:16) and four activation transformation methods: S-PTS, D-PTS, R-Sparse, and VAR.
-First of all, we observe a strong model degradation on generative tasks.
-Second of all, we see that VAR is the strongest performer overall, especially for Llama3-8B.
-S-PTS/D-PTS are competitive and lightweight, and R-Sparse lags significantly, particularly at 2:4.
-We speculate that while semi-structured patterns are good for prefill stage in LLMs they significantly degrade performance during decode stage.
-However, as discussed in Section 1 decode stage for single vector can be accelerated with more flexible approaches.
-Table 3: Evaluation with Instruction-Following (IFEval) for Llama3-8B and Qwen2.5-7B. PS denotes prompt-level strict acc, PL denotes prompt-level loose acc.
-Model Original Pattern S-PTS D-PTS R-Sparse VAR
-PS PL PS PL PS PL PS PL PS PL
-Llama3-8B 0.4455 0.4861 2:4 0.1682 0.1904 0.1941 0.2015 0.0869 0.0979 0.2237 0.2458
-8:16 0.2995 0.3327 0.2828 0.3198 0.2089 0.2311 0.3161 0.3586
-Qwen2.5-7B 0.7135 0.7394 2:4 0.4325 0.5176 0.4399 0.5120 0.2736 0.3457 0.4565 0.5342
-8:16 0.5194 0.5804 0.5434 0.5989 0.3697 0.4196 0.5249 0.5896
-4.6 Selection Criteria and Transformations for Unstructured Pruning
-Although unstructured pruning offers limited efficiency gains compared to structured variants, it serves as a valuable proxy task for evaluating the robustness and generalization of proposed methods.
-We evaluate D-PTS, VAR, and two selection criteria in this experiment: CLACT and Amber-Pruner using the Llama3-8B model.
-Results in Table 4 indicate that VAR is the most effective transformation under unstructured sparsity.
-Moreover, CLACT outperforms Amber-Pruner by a wider margin here than in our semi-structured pruning experiments.
-These findings suggest two key insights: (1) No single method emerges as optimal for both unstructured and semi-structured sparsity.
-(2) The methods proposed in this work, VAR and CLACT, demonstrate strong generalization and are well-suited for unstructured activation pruning.
-Table 4: Performance comparison of pruning methods under 50% and 70% unstructured sparsity for Llama3-8B model.
-Model ARC Easy BoolQ PIQA WinoGrande Avg. Drop (%)
-Original 0.8207 0.8391 0.8003 0.7340
-Unstructured 50%
-ACT 0.7770 0.8198 0.7714 0.6858 4.45
-D-PTS 0.78577 0.8253 0.7807 0.68981 3.60
-VAR 0.78409 0.8192 0.77584 0.7048 3.47
-CLACT 0.7803 0.8253 0.7655 0.70007 3.89
-Amber-Pruner 0.768 0.82018 0.7627 0.7016 4.45
-Unstructured 70%
-ACT 0.55808 0.63119 0.6474 0.5477 25.32
-D-PTS 0.56481 0.62415 0.6507 0.53433 25.68
-VAR 0.61447 0.65107 0.67573 0.5319 22.66
-CLACT 0.5551 0.6039 0.62676 0.52407 27.67
-Amber-Pruner 0.48737 0.5938 0.5897 0.539 30.68
-4.7 Combination of Transformations and Pruning criteria
-Next, we evaluated combinations of multiple approaches to explore potential performance gains.
-These combinations and their results are presented in Table 7.
-As shown, none of the evaluated combinations outperforms any single method, highlighting the challenges of naively combining them.
-Table 5: Aggregated results on Llama3-8B with 8:16 activation sparsity with different layers pruned.
-ORIG.
-AVG.
-denotes original model performance.
-Layers indicates the subset of linear layers where the method was applied.
-Non aggregated results are presented in Appendix 12.
-Drop is computed without accounting for perplexity.
-ORIG. AVG. Method Layers PPL AVG. Drop $\downarrow$
-LS+L-PTS all 9.6036 0.6047 10.90%
-LS+L-PTS key,out,gate,down 8.3483 0.6385 5.43%
-0.6726 LS+L-PTS key,value,gate,down 8.0821 0.6503 3.56%
-LS+L-PTS + VAR all 9.4983 0.6056 10.60%
-LS+L-PTS + VAR key,out,gate,down 8.2930 0.6422 4.64%
-LS+L-PTS + VAR key,value,gate,down 8.0259 0.6516 3.36%
-4.8 Layer Sensitivity
-The next question we investigate in this work is layer sensitivity to activation sparsity.
-For this experiment, we use the Extended Datasets, the Llama3-8B model, and the 8:16 sparsity pattern.
-The main results are presented in Table 5.
-Here, we focus on methods with learnable parameters, primarily L-PTS and LS (a learnable, diagonal scaling projection).
-While in most of our experiments, learnable methods underperformed on average compared to static or magnitude-based approaches, the best-performing configurations for Llama3-8B under 8:16 sparsity were, in fact, those employing learnable parameters — as shown in Table 5.
-We find that activations from the up projection (in the FFN) and the out projection [^1] The output projection of the attention mechanism.
-It combines outputs from all attention heads and projects them back to the model’s hidden dimension.
-are the most sensitive to sparsification, pruning these layers leads to the most significant performance drops, suggesting they should be preserved or treated with greater care in compression strategies.
-While we can not say this generalizes across all layers, we empirically demonstrate that some layers are more important than others.
-5 Discussion
-The performance gap between multiple/single-choice benchmarks (e.g., BoolQ, PIQA) and IFEval likely stems from differences in the inference stages they emphasize.
-Core QA benchmarks primarily stress the prefill phase, whereas IFEval evaluates both prefill and autoregressive generation.
-Our evaluation remains valid: semi-structured patterns like 2:4 and 8:16 are especially effective at accelerating the prefill stage, which often dominates inference latency.
-6 Limitations
-Our work has three main limitations.
-First, evaluations rely on software emulation, precluding real-world measurements of speedup or energy savings on hardware supporting sparsity beyond 2:4.
-Second, our layer sensitivity analysis is preliminary, while up-projection and out-projection layers appear most vulnerable, broader architectural studies are needed.
-Third, Qwen2.5-7B shows anomalous gains on multiple-choice benchmarks under certain sparsification methods, an artifact we attribute to benchmark limitations, as no such gains appear on IFEval.
-Critically, generative performance degrades significantly, underscoring the need for evaluation beyond multiple-choice QA.
-7 Conclusion
-This work establishes that post-training activation pruning is more accuracy preserving than weight pruning for Large Language Models.
-Across four diverse LLMs, we demonstrate that activation pruning consistently preserves a model’s capabilities better than weight pruning at equivalent sparsity levels, highlighting its potential for dynamic, input-adaptive efficiency gains.
-Our key novel contributions are twofold.
-First, we conduct the first comprehensive evaluation of lightweight, plug-and-play error mitigation techniques for activation sparsity, including the introduction and evaluation of three new methods: Cosine Loss Activation (CLACT) as a context-aware pruning criterion, Dynamic/Learnable Per-Token Shift (D-PTS/L-PTS) and Variance Correction (VAR).
-We show that these simple methods, often outperforming more complex approaches, providing strong, hardware-friendly baselines for the community.
-Second, we systematically explore semi-structured sparsity patterns beyond the hardware standard 2:4.
-We find that the 8:16 pattern offers more flexibility and, ultimately, lower model degradation.
-Although the 16:32 pattern performs even closer to unstructured sparsity, we advocate for 8:16 as the optimal target for future hardware design due to its balance of performance gain and implementation feasibility.
-In summary, this work not only provides practical, high-performing methods for post-training activation pruning but also creates grounded motivation for 8:16 structured sparsity support, which would unlock significant efficiency gains for LLM inference with minimal added complexity.
-Acknowledgment on LLM assisted writing:
-This paper used open access Qwen3-Max, in some parts of the paper, for proofreading and text rephrasing in accordance with formal style.
-Appendix
-Appendix A R-Sparse Details
+Appendix B R-Sparse Details
 Finally, we include R-Sparse (Kamirul et al., 2025), which combines activation sparsity with a low-rank approximation of the weight matrix.
@@
 where
-$\mathbf{Y}_{s}=\sigma_{t(s)}(\mathbf{X})\mathbf{W}^{\top},\quad\mathbf{Y}_{r}=(\mathbf{X}-\sigma_{t(s)}(\mathbf{X}))(A_{r}B_{r})^{\top}.$ (8)
+$\mathbf{Y}_{s}=\sigma_{t(s)}(\mathbf{X})\mathbf{W}^{\top},$
+$\mathbf{Y}_{r}=(\mathbf{X}-\sigma_{t(s)}(\mathbf{X}))(A_{r}B_{r})^{\top}.$
 Here $\sigma_{t(s)}(\cdot)$ denotes sparsification of activations with threshold $t(s)$, and $A_{r}B_{r}^{\top}$ is the rank-$r$ approximation of $\mathbf{W}$ obtained from its truncated SVD.
 The trade-off between $\mathbf{Y}_{s}$ and $\mathbf{Y}_{r}$ is determined by a sparsity budget $s$ and rank $r$, which can be optimized via evolutionary search.
-Appendix B Different patterns for semi-structure sparsification
-Table 6: Performance comparison of different sparsity patterns on Llama3-8B across various benchmarks.
+Appendix C Main extended results
+Here, we present un-aggregated results.
+Comparisons between different sparsity patterns is presented in Table 7.
+In Table 11 and Table 12 we present the results of the error mitigation strategies and the selection criteria.
+Table 11 reports the results of the combined methods.
+Table 10 compares unstructured and semi-structured sparsity.
+Finally, in Table 13 we demonstrate results when some of the layers are excluded to evaluate layer sensitivity.
+Table 7: Performance comparison of different sparsity patterns on Llama3.1-8B-Intsruct across various benchmarks.
 Values represent accuracy scores, with the last column showing the average performance drop relative to the original model.
@@
 70% unstructured $0.5580$ $0.6311$ $0.6474$ $0.5477$ 25.3225.32%
-B.1 Comparison of different patterns
 Table 8: A comparison of combined approaches with 8:16 semi-structured sparsity.
@@
 Models
-Method Llama2-7B Qwen2.5-7B Gemma3-4B LLama3-8B Average Drop ($\downarrow)$
+Method Llama2-7B-chat Qwen2.5-7B-Instruct Gemma3-4B-Instruct Llama3.1-8B-Instruct Average Drop ($\downarrow)$
 CLACT + PTS 5.635.63% 5.06-5.06% 0.500.50% 8.558.55% 2.40%
@@
 L-PTS + VAR 6.876.87% 2.862.86% 3.413.41% 7.157.15% 5.075.07%
-Appendix C Datasets
+Appendix D Datasets
+Detailed description of the datasets is given in Table 9.
 Table 9: Datasets used to evaluate hypotheses.
@@
 PIQA (Bisk et al., 2020) Physical commonsense QA benchmark for choosing the right answer between two options. Contains 16K train, 2K dev, and 3K test examples. Accuracy
-WinoGrande (Sakaguchi et al., 2020) QA benchmark for pronoun resolution with adversarial filtering. Contains 40K train, 1267 dev, and 1767 test examples. Accuracy
+WinoGrande (Sakaguchi et al., 2021) QA benchmark for pronoun resolution with adversarial filtering. Contains 40K train, 1267 dev, and 1767 test examples. Accuracy
 HellaSwag (Zellers et al., 2019) Commonsense reasoning benchmark for sentence completion, designed to be easy for humans but hard for models. Contains 70K train and 10K validation examples. Accuracy
 OpenBookQA (Mihaylov et al., 2018) Open-book question answering dataset requiring retrieval of elementary science facts. Contains 5957 4-way multiple-choice questions. Accuracy
-RTE (Dagan et al., 2005; Bar-Haim et al., 2006) Recognizing Textual Entailment datasets from PASCAL challenges. Task is to classify if a hypothesis is entailed by a premise. Accuracy
+RTE (Dagan et al., 2005; Haim et al., 2006) Recognizing Textual Entailment datasets from PASCAL challenges. Task is to classify if a hypothesis is entailed by a premise. Accuracy
 MMLU (Hendrycks et al., 2020) Massive Multitask Language Understanding benchmark covering 57 subjects across STEM, humanities, and social sciences. Measures multitask accuracy. Accuracy
@@
 IFEval (Zhou et al., 2023) Benchmark with 541 prompts containing verifiable instructions to measure instruction-following fidelity. Accuracy (Prompt-level) Accuracy (Instruct-level)
-Appendix D Weights versus Activations
 Table 10: The performance of models with applied unstructured activation pruning.
@@
 ACT stands for activations pruning, WT — for weight pruning.
-OUT denotes values more than $2^{3}$, according accuracy scores most likely correspond to random.
+OUT denotes values more than $10^{3}$, according accuracy scores most likely correspond to random.
 Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Drop $(\downarrow)\%$
-Llama2-7B
+Llama2-7B-chat
    Base 6.94 0.74 0.80 0.76 0.66 -
@@
    0.9 WT OUT 0.27 0.38 0.53 0.48 $43.39\%$
-Qwen2.5-7B
+Qwen2.5-7B-Instruct
    Base 7.46 0.69 0.86 0.75 0.60 -
@@
    0.9 WT OUT 0.25 0.58 0.54 0.51 $\textbf{36.35}\%$
-Gemma3-4B
+Gemma3-4B-Instruct
    Base 17.29 0.72 0.84 0.72 0.62 -
@@
    0.9 WT OUT 0.25 0.45 0.52 0.52 $\textbf{40.60}\%$
-Appendix E Main extended results
-In this section we present non - aggregated results for all of the methods and datasets., results are presented in Table 11 and Table 10.
-Table 10: Semi-Structured 2:4 Sparsification - performance Metrics, for calibration, when it is required, and perplexity we use WikiText2. Average Drop is computed without accounting for perplexity
+Table 11: Semi-Structured 2:4 Sparsification - performance Metrics, for calibration, when it is required, and perplexity we use WikiText2. Average Drop is computed without accounting for perplexity.
 Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Average Drop %
-   Llama2-7B 6.94 0.74 0.80 0.76 0.66 -
+   Llama2-7B-chat 6.94 0.74 0.80 0.76 0.66 -
 ACT 10.23 0.66 0.71 0.71 0.60 9.43%
@@
 R-SPARSE (128) 9.29 0.65 0.65 0.70 0.59 12.23%
-   Llama3.1-8B 7.21 0.82 0.84 0.80 0.73 -
+   Llama3.1-8B-Instruct 7.21 0.82 0.84 0.80 0.73 -
 ACT 16.61 0.68 0.73 0.72 0.61 14.35%
@@
 R-SPARSE (128) 16.09 0.67 0.71 0.70 0.61 16.34%
-   Qwen2.5-7B 7.46 0.69 0.86 0.75 0.60 -
+   Qwen2.5-7B-Instruct 7.46 0.69 0.86 0.75 0.60 -
 ACT 10.06 0.65 0.86 0.72 0.54 4.95%
@@
 R-SPARSE (128) 9.12 0.77 0.77 0.75 0.63 -1.51%
-   Gemma3-4B 17.29 0.72 0.84 0.72 0.62 -
+   Gemma3-4B-Instruct 17.29 0.72 0.84 0.72 0.62 -
 ACT 35.62 0.65 0.76 0.70 0.51 9.94%
@@
 Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Average Drop %
-   Llama2-7B 6.94 0.74 0.80 0.76 0.66 -
+   Llama2-7B-chat 6.94 0.74 0.80 0.76 0.66 -
 ACT 8.12 0.69 0.75 0.73 0.63 5.37%
@@
 LPTS + VAR 7.52 0.69 0.67 0.74 0.64 6.87%
-   Llama3.1-8B 7.21 0.82 0.84 0.80 0.73 -
+   Llama3.1-8B-Instruct 7.21 0.82 0.84 0.80 0.73 -
 ACT 10.32 0.75 0.80 0.76 0.66 7.38%
@@
 LPTS + VAR 10.26 0.77 0.78 0.76 0.66 7.15%
-   Qwen2.5-7B 7.46 0.69 0.86 0.75 0.60 -
+   Qwen2.5-7B-Instruct 7.46 0.69 0.86 0.75 0.60 -
 ACT 8.61 0.66 0.87 0.73 0.53 4.38%
@@
 LPTS + VAR 8.21 0.70 0.83 0.73 0.56 2.80%
-   Gemma3-4B 17.29 0.72 0.84 0.72 0.62 -
+   Gemma3-4B-Instruct 17.29 0.72 0.84 0.72 0.62 -
 ACT 25.31 0.70 0.81 0.71 0.55 4.76%
@@
 LPTS + VAR 15.68 0.71 0.79 0.72 0.57 3.41%
-Table 12: Llama3-8B with 8:16 activation sparsity.
+Table 13: Llama3.1-8B-Instruct with 8:16 activation sparsity.
 LS+L-PTS indicates Learnable Diagonal Scale + Learnable Shift, “Layers” indicates the subset of linear layers where the method was applied.